### v2

# Q-learning with Posterior Sampling

Abstract
Bayesian posterior sampling techniques have demonstrated superior empirical performance in many exploration-exploitation settings. However, their theoretical analysis remains a challenge, especially in complex settings like reinforcement learning. In this paper, we introduce Q-Learning with Posterior Sampling (PSQL), a simple Q-learning-based algorithm that uses Gaussian posteriors on Q-values for exploration, akin to the popular Thompson Sampling algorithm in the multi-armed bandit setting. We show that in the tabular episodic MDP setting, PSQL achieves a regret bound of $\tilde{O}(H^{2}\sqrt{SAT})$, closely matching the known lower bound of $\Omega(H\sqrt{SAT})$. Here, S, A denote the number of states and actions in the underlying Markov Decision Process (MDP), and $T=KH$ with $K$ being the number of episodes and $H$ being the planning horizon. Our work provides several new technical insights into the core challenges in combining posterior sampling with dynamic programming and TD-learning-based RL algorithms, along with novel ideas for resolving those difficulties. We hope this will form a starting point for analyzing this efficient and important algorithmic technique in even more complex RL settings.
1 Introduction
In a Reinforcement Learning (RL) problem, an agent interacts sequentially with an unknown environment and uses the observed outcomes to learn an interaction strategy. The underlying mathematical model for RL is a Markov Decision Process (MDP). In the tabular episodic RL setting, there is an underlying MDP with a finite state space $\cal S$}, a finite action space $\cal A$} and a planning horizon $H$. On taking an action $a$ in state $s$ at step $h$, the environment produces a reward and next state from the (unknown) reward model $R_{h}(s,a)$ and transition probability model $P_{h}(s,a)$ of the underlying MDP.
Q-learning Watkins and Dayan (1992) is a classic dynamic programming (DP)-based algorithm for RL. The DP equations (aka Bellman equations) provide a recursive expression for the optimal expected reward achievable from any state and action of the MDP, aka the $Q$-values, in terms of the optimal value achievable in the next state. Specifically, for any given $s\in{\cal S},a\in{\cal A},h\in[H]$, the Q-value $Q_{h}(s,a)$ is given by:
$Q_{h}(s,a)=\max_{a\in{\cal A}}R_{h}(s,a)+\sum_{s^{\prime}\in{\cal S}}P_{h}(s,a,s^{\prime})V_{h+1}(s^{\prime}),\text{ with}$ (1)
$V_{h+1}(s^{\prime})\mathrel{\mathop{\ordinarycolon}}=\max_{a^{\prime}\in{\cal A}}Q_{h+1}(s^{\prime},a^{\prime}).$
Here, the terminal value $V_{H+1}(s)$ is 0 for all $s$. The optimal action in state $s$ is then given by the argmax action in the above.
In the reinforcement learning setting, where the reward and transition models of the MDP are unknown, the $Q$-learning algorithm uses the celebrated Temporal Difference (TD) learning idea Sutton (1988) to construct increasingly accurate estimates of $Q$-values using past observations. The key idea here is to construct an estimate of the right hand side of the Bellman equation, aka target, by bootstrapping the current estimate $\widehat{V}_{h+1}$ for the next step value function. That is, on playing an action $a$ in state $s$ at step $h$, and observing reward $r_{h}$ and next state $s^{\prime}$, the target $z$ is typically constructed as:
$z\mathrel{\mathop{\ordinarycolon}}=r_{h}+\widehat{V}_{h+1}(s^{\prime})$
And then the estimate ${\widehat{Q}}_{h}(s,a)$ for $Q_{h}(s,a)$ is updated to fit the Bellman equations using the Q-learning update rule [^2] Or, $\widehat{Q}_{h}(s,a)\leftarrow\widehat{Q}_{h}(s,a)+\alpha_{n}(z-\widehat{Q}_{h}(s,a))$ where $z-\widehat{Q}_{h}(s,a)$ is called the Temporal Difference (TD).
$\widehat{Q}_{h}(s,a)\leftarrow(1-\alpha_{n})\widehat{Q}_{h}(s,a)+\alpha_{n}z.$ (2)
Here $\alpha_{n}$ is an important parameter of the $Q$-learning algorithm, referred to as the learning rate. It is typically a function of the number of previous visits $n$ for the state $s$ and action $a$.
There are several ways to interpret the Q-learning update rule. The traditional frequentist interpretation popularized by Mnih et al. (2015) interprets this update as a gradient descent step for a least squares regression error minimization problem. Another, arguably more insightful, interpretation of $Q$-learning can be obtained using Bayesian inference theory. Specifically, if we assume a Gaussian prior
${\cal N}(\widehat{Q}_{h}(s,a),\frac{\sigma^{2}}{n-1})$
on the $Q$-value $Q_{h}(s,a)$, and a Gaussian likelihood function ${\cal N}(Q_{h}(s,a),\sigma^{2})$ for the target $z$, then using Bayes rule, one can derive the Bayesian posterior as the Gaussian distribution
${\cal N}(\widehat{Q}_{h}(s,a),\frac{\sigma^{2}}{n}),\qquad\qquad\text{ where }\widehat{Q}_{h}(s,a)\leftarrow(1-\alpha_{n})\widehat{Q}_{h}(s,a)+\alpha_{n}z$ (3)
(See Section 3.1 for further details.)
Importantly, the Bayesian posterior tracks not just the mean but also the variance or uncertainty in the $Q$-value estimate. This uncertainty quantification is particularly useful for managing the exploration-exploitation tradeoff for regret minimization in reinforcement learning.
Figure 1: A comparison of PSQL(a heuristic derived from PSQL), UCBQL Jin et al. (2018); (Staged) RandQLTiapkin et al. (2023), and RLSVI Russo (2019) in a chain MDP environment (for details, and more experiments, see Appendix A).
Intuitively, the state and actions with a small number of past visits (i.e., small $n$) have large uncertainty in their current $Q$-value estimate, and should be explored more. The posterior sampling approaches implement this idea by simply taking a sample from the posterior, which is likely to be closer to the mean (less exploration) for actions with small posterior variance, and away from the mean (more exploration) for those with large variance.
Following this intuition, we introduce Q-learning with Posterior Sampling (PSQL) algorithm that maintains a posterior on Q-values for every state and action. Then, to decide an action in any given state, it simply generates a sample from the posterior for each action, and plays the arg max action of the sampled $Q$-values.
Popularized by their success in the multi-armed bandit settings Thompson (1933); Chapelle and Li (2011); Kaufmann et al. (2012); Agrawal and Goyal (2017), and in deep reinforcement learning regimes (Osband et al., 2016a; Fortunato et al., 2017; Azizzadenesheli et al., 2018; Li et al., 2021b; Fan and Ming, 2021; Sasso et al., 2023), the posterior sampling approaches are generally believed to be more efficient in managing the exploration-exploitation tradeoff than their UCB (Upper Confidence Bound) counterparts. Our preliminary experiments (see Figure 1) suggest that this is also the case for our Q-learning approach in the tabular RL setting.  [^3] For the empirical study reported here, we implement a vanilla version of posterior sampling with Q-learning. The PSQL algorithm presented later modifies the target computation slightly as described later in Section 3, for the sake of theoretical analysis. However, obtaining provable guarantees for posterior sampling approaches have historically been more challenging.
Several previous works (e.g., Li et al. (2021a); Jin et al. (2018)) use UCB-based exploration bonuses to design optimistic Q-learning algorithms with near-optimal regret bounds. [^4] The algorithm from Jin et al. (2018) is referred as UCBQL in the text and experiments. For the posterior sampling based approaches however, the first tractable $Q$-learning based algorithm with provable regret bounds was provided only recently by Tiapkin et al. (2023) for the Staged-RandQL algorithm. However, the Staged-RandQL (and RandQL) algorithm presented in their work deviated from the natural approach of putting a posterior on Q-values, and instead derived a Dirichlet Bayesian posterior on the transition probabilities, which is conceptually closer to some model-based posterior sampling algorithms, e.g., the PSRL algorithm in Agrawal and Jia (2017). RandQL implements sampling from the implied distribution on Q-estimates in a more efficient way via learning rate randomization, so that it qualifies as a model-free algorithm.
Another closely related approach with provable regret bounds is the RLSVI (Randomized Least Squared Value Iteration) algorithm by Osband et al. (2016b, 2019); Russo (2019); Zanette et al. (2020); Agrawal et al. (2021); Xiong et al. (2022). The RLSVI algorithm is an approximate value iteration-based approach that can be interpreted as maintaining “empirical posteriors" over the value functions by injecting noise. However, in the tabular setting considered here, RLSVI reduces to a model-based algorithm. (See Section 2.1 for further comparison.) There are several other model-based posterior sampling algorithms in the literature with near optimal regret bounds Osband et al. (2013); Osband and Van Roy (2017); Ouyang et al. (2017); Agrawal and Jia (2022); Agarwal and Zhang (2022); Tiapkin et al. (2022). Model-based algorithms directly estimate the reward and/or transition model, instead of the implied optimal value functions, or policy parameters. In many settings, model-based algorithms can be more sample efficient. But, model-free approaches like Q-learning have gained popularity in practice because of their simplicity and flexibility, and underlie most successful modern deep RL algorithms (e.g., DQN (Mnih et al., 2013), DDQN (van Hasselt et al., 2015), A3C (Mnih et al., 2016)). Provable regret bounds for a simple posterior sampling based Q-learning algorithm like PSQL, therefore, still remains a problem of significant interest.
Our contributions are summarized as follows.
  • We propose the Q-learning with Posterior Sampling (PSQL) algorithm that combines the simplicity and flexibility of the popular Q-learning algorithm with efficient exploration provided by the Bayesian posterior sampling approach. Our preliminary experiments demonstrate promising empirical performance of this simple algorithm compared to contemporary approaches. (See Section 3 for algorithm design and Appendix A for experiments.)
  • We provide a novel derivation of Q-learning as a solution to a Bayesian inference problem with a regularized Evidence Lower Bound (ELBO) objective. Besides forming the basis of our PSQL algorithm design, this derivation provides a more insightful interpretation of the modified learning rates introduced in some previous works on $Q$-learning (e.g., (Jin et al., 2018)) to obtain provable regret bounds. (See Section 3.1.)
  • We prove a near-optimal regret bound of $\tilde{O}(H^{2}\sqrt{SAT})$ for PSQL which closely match the known lower bound of $\Omega(H\sqrt{SAT})$ (Jin et al., 2018). Our result improves the regret bounds available for the closely related approach of RLSVI Russo (2019) and matches those recently derived by Tiapkin et al. (2023) for a more complex posterior sampling based algorithm Staged-RandQL. (See Section 2,4.)
  • Our regret analysis reveals several key difficulties in combining posterior sampling with DP and TD-learning-based algorithms due to error accumulation in the bootstrapped target; along with novel ideas for overcoming these challenges. (See Section 4.1.)
2 Our setting and main result
We consider an online reinforcement learning setting. In this setting, the algorithm interacts with environment in $K$ sequential episodes, each containing $H$ steps. At step $h=1,\ldots,H$ of each episode $k$, the algorithm observes the current state $s_{k,h}$, takes an action $a_{k,h}$ and observes a reward $r_{k,h}$ and the next state $s_{k,h+1}$.
The reward and next state are generated by the environment according to a fixed underlying MDP $({\cal S},{\cal A},R,P)$, so that $\Pr(s_{k,h+1}=s^{\prime}|s_{k,h}=s,a_{k,h}=a)=P_{h}(s,a,s^{\prime}),\mathbb{E}[r_{k,h}|s_{k,h}=s,a_{k,h}=a]=R_{h}(s,a)$. However, the reward functions and the transition probability distributions $R_{h},P_{h},h=1,\ldots,H$ are apriori unknown to the algorithm. The goal is to minimize total regret compared to the optimal value given by the dynamic programming equation (1). Specifically, let $\pi_{k}$ denote the policy used by the algorithm in episode $k$, so that $a_{k,h}=\pi^{k}(s_{k,h})$. We aim to bound regret, defined as
$\text{Reg}(K)\mathrel{\mathop{\ordinarycolon}}=\sum_{k=1}^{K}(V_{1}(s_{k,1})-V^{\pi_{k}}(s_{k,1}))$ (4)
for any set of starting states $s_{k,1},k=1,\ldots,K$.
A regret minimizing algorithm uses past observations to learn about the MDP and improve its actions. Since the algorithm can only observe the environment’s response at a visited state and action, a main challenge in this problem is managing the exploration-exploitation tradeoff. This refers to the dilemma between picking actions that are most likely to be optimal according to the observations made so far, versus picking actions for that allow visiting less-explored states and actions. Following work on managing exploration-exploitation tradeoffs in multi-armed bandit (MAB) settings, two main approaches for exploration exploitation are optimistic approaches based on UCB and posterior sampling approaches (aka Thompson Sampling in MAB).
[…]
Theorem 1 (Informal).
The cumulative regret of our PSQL (Algorithm 1 ) in $K$ episodes with horizon $H$ is bounded as $\text{Reg}(K)\leq\tilde{O}\mathinner{\left(H^{2}\sqrt{SAT}\right)},$ where $T=KH$.
Next, we derive the posterior sampling approach considered in this paper, and relate it to the other popular versions of Q-learning for regret minimization in RL.
2.1 Related works
Algorithm Regret Comments
UCB UCBQL Jin et al. (2018) $\tilde{O}(H^{1.5}\sqrt{SAT})$ Q-learning with UCB
Q-EarlySettled-Advantage Li et al. (2021a) $\tilde{O}(H\sqrt{SAT})$ Q-learning with UCB
Posterior sampling Conditional Posterior Sampling Dann et al. (2021) $\tilde{O}(HSA\sqrt{T})$ computationally intractable
RLSVI Russo (2019) $\tilde{O}(H^{3}S^{1.5}\sqrt{AT})$ approximate value iteration with noise
C-RLSVI Agrawal et al. (2021) $\tilde{O}(H^{2}S\sqrt{AT})$ approximate value iteration with noise
Staged RandQL Tiapkin et al. (2023) $\tilde{O}(H^{2}\sqrt{SAT})$ Q-learning with learning-rate randomization
PSQL [this work] $\tilde{O}(H^{2}\sqrt{SAT})$ Q-learning with Gaussian posteriors
Lower bound Jin et al. (2018) $\Omega(H\sqrt{SAT})$ -
Table 1: Comparison of our regret bound to related wors ( Dann et al. (2021) is in function approximation setting).
Our work falls under the umbrella of online episodic reinforcement learning, specifically on regret minimization in tabular setting. In the category of the Upper Confidence Bound (UCB)-based algorithms, there is a huge body of research both on model-based (Bartlett and Tewari, 2012; Azar et al., 2017; Fruit et al., 2018; Zanette and Brunskill, 2019; Zhang et al., 2020; Boone and Zhang, 2024), and model-free Jin et al. (2018); Bai et al. (2019); Ménard et al. (2021); Zhang and Xie (2023); Agrawal and Agrawal (2024) algorithms. In-fact, Jin et al. (2018) were the first to provide a near-optimal worst-case regret bound of $\tilde{O}(\sqrt{H^{3}SAT})$, subsequently improved to $\tilde{O}(H\sqrt{SAT})$ by Zhang et al. (2020); Li et al. (2021a).
Motivated by the superior empirical performance of Bayesian posterior sampling approaches compared to their UCB counterparts (Chapelle and Li, 2011; Kaufmann et al., 2012; Osband et al., 2013; Osband and Van Roy, 2017; Osband et al., 2019) there have been several attempts at deriving provable regret bounds for these approaches in the episodic RL setting. Among model-based approaches, near optimal regret bounds have been established for approaches that use (typically Dirichlet) posteriors on transition models Ouyang et al. (2017); Agrawal and Jia (2017); Tiapkin et al. (2022). There have been relatively limited studies on model-free, sample-efficient and computationally efficient Bayesian algorithms. Dann et al. (2021) proposed one such framework but is computationally intractable. Our work aims to fill this gap.
A popular approach closely related to posterior sampling is Randomized Least Square Value Iteration (RLSVI) (Osband et al., 2016b, 2019; Russo, 2019; Zanette et al., 2020). In RLSVI, the exploration is carried out by injecting randomized uncorrelated noise to the reward samples, followed by a re-fitting of a $Q$-function estimate by solving a least squares problem on all the past data, incurring heavy computation and storage costs. This process has been interpreted as forming an approximate posterior distribution over value functions. RLSVI too enjoys a “superior-than-UCB” empirical performance. However, its worst-case regret bounds Russo (2019) remain suboptimal (see Table 1) in their dependence on the size of the state space.
More recently, Tiapkin et al. (2023) proposed (RandQL and Staged-RandQL) algorithms that are model-free, tractable and enjoy $\tilde{O}(H^{2}\sqrt{SAT})$ regret by randomizing the learning rates of the Q-learning update rule. Their algorithmic design is based on Dirichlet posteriors on transition models and efficient implmentaton of the implied distribution on $Q$-value estimates via learning rate randomization. In our preliminary experiments (see Figure 1 and Section A), our PSQL approach with simple Gaussian based posteriors shows significantly better performance compared to these algorithms.
In Table 1 we provide a detailed comparison of our results with the above-mentioned related work on posterior sampling algorithms for RL.
3 Algorithm design
We present our Q-learning with Posterior Sampling (PSQL) algorithm. We first present a Bayesian posterior-based derivation of the Q-learning update rule, which forms the basis for our algorithm design.
3.1 Posterior derivation
An insightful interpretation of the Q-learning update rule can be obtained using Bayesian inference. Let $\theta$ denote the Bayesian parameter that we are inferring, which in our case is the quantity $Q_{h}(s,a)$. Given a prior $p$ on $\theta$, log likelihood function $\ell(\theta,\cdot)$, and a sample $z$, the Bayesian posterior $q$ is given by the Bayes rule:
$q(\theta)\propto p(\theta)\cdot\exp(\ell(\theta,z))$ (5)
which can also be derived as an optimal solution of the following optimization problem (see Chapter 10 in Bishop and Nasrabadi (2006)), whose objective is commonly referred to as Evidence Lower Bound (ELBO):
$\max_{q}\mathbb{E}_{\theta\sim q}[\ell(\theta,z)]-KL(q||p)$ (6)
where $KL(\cdot||\cdot)=\int_{\theta}q(\theta)\log(\frac{q(\theta)}{p(\theta)})$ denotes KL-divergence function. It is well known that when $p(\cdot)$ is Gaussian, say ${\cal N}(\widehat{\mu},\frac{\sigma^{2}}{n-1})$, and the likelihood given $\theta$ is Gaussian ${\cal N}(\theta,\sigma^{2})$, then the posterior $q(\cdot)$ is given by the Gaussian distribution
${\cal N}(\widehat{\mu},\frac{\sigma^{2}}{n}),\text{ with }\widehat{\mu}\leftarrow(1-\alpha_{n})\widehat{\mu}+\alpha_{n}z$ (7)
with $\alpha_{n}=\frac{1}{n}$. Therefore, substituting $\theta$ as $Q_{h}(s,a)$ and $\widehat{\mu}$ as $\widehat{Q}_{h}(s,a)$, the above yields the Q-learning learning update rule (2) with learning rate $\alpha_{n}=\frac{1}{n}$.
A caveat is that the above assumes $z$ to be an unbiased sample from the target distribution, whereas in Q-learning, $z$ is biased due to bootstrapping. In a recent work, Jin et al. (2018) observed that in order to account for this bias and obtain theoretical guarantees for Q-learning, the learning rate needs to be adjusted to $\alpha_{n}=\frac{H+1}{H+n}$. In fact, Bayesian inference can also provide a meaningful interpretation of this modified learning rate proposed in Jin et al. (2018). Consider the following “regularized" Bayesian inference problem Khan and Rue (2023) which adds an entropy term to the ELBO objective in (6):
$\max_{q}\mathbb{E}_{\theta\sim q}[\ell(\theta,z)]-KL(q||p)+\lambda_{n}{\cal H}(q)$ (8)
where ${\cal H}(q)$ denotes the entropy of the posterior. Then, we show (Lemma 7) that with weight $\lambda_{n}\mathrel{\mathop{\ordinarycolon}}=\frac{H}{n}$, when the prior $p(\theta)$ is Gaussian ${\cal N}(\widehat{\mu},\frac{\sigma^{2}}{n-1})$, and the likelihood of $z$ given $\theta$ is Gaussian ${\cal N}(\theta,\frac{\sigma^{2}}{H+1})$, then the posterior $q(\cdot)$ is given by the Gaussian distribution in (7), but with the learning rate $\alpha_{n}=\frac{H+1}{H+n}$ suggested in Jin et al. (2018). A derivation is provided in Appendix B.
Intuitively, since we suspect bias in the target $z$, (8) introduces additional uncertainty in the posterior via entropy. The weight $\lambda_{n}$ of this entropy term decreases as the number of samples increases and the bootstrapped target is expected to have lower bias. This derivation may be of independent interest as it provides a more intuitive explanation of the modified learning rate schedule proposed in Jin et al. (2018), where it was motivated mainly by the mechanics of regret analysis.
Substituting $\theta$ as $Q_{h}(s,a)$ and $\widehat{\mu}$ has $\widehat{Q}_{h}(s,a)$ in (8), we derive that given a Gaussian prior ${\cal N}(\widehat{Q}_{h}(s,a),\frac{\sigma^{2}}{n-1})$ over $Q_{h}(s,a)$, Gaussian likelihood ${\cal N}(Q_{h}(s,a),\frac{\sigma^{2}}{H+1})$ on target $z$, the following posterior maximizes the regularized ELBO objective:
${\cal N}(\widehat{Q}_{h}(s,a),\frac{\sigma^{2}}{n}),\,\text{where}\qquad\widehat{Q}_{h}(s,a)\leftarrow(1-\alpha_{n})\widehat{Q}_{h}(s,a)+\alpha_{n}z,$ (9)
where $\alpha_{n}=\frac{H+1}{H+n}$, and $n$ is the number of samples for $s,a$ observed so far.
The above posterior derivation forms the basis of our algorithm design presented next.
3.2 Algorithm details
A detailed pseudo-code of our PSQL algorithm is provided as Algorithm 1,2. It uses the current Bayesian posterior to generate samples of $Q$- values at the current state and all actions, and plays the arg max action. Specifically, at a given episode $k$, let $s_{h}$ be the current state observed in the beginning of the episode and for action $a$, let $N_{h}(s_{h},a)$ be the number of visits of state $s_{h}$ and action $a$ before this episode. Let $\widehat{Q}_{h}(s_{h},a)$ be the current estimate of the posterior mean, and
$\textstyle\sigma(n)=\frac{\sigma^{2}}{n+1}\mathrel{\mathop{\ordinarycolon}}=64\frac{H^{3}}{n+1}\log(1/\delta).$ (10)
Then, the algorithm samples for each $a$,
$\tilde{Q}_{h}(s_{h},a)\sim{\cal N}(\widehat{Q}_{h}(s_{h},a),\sigma(N_{h}(s_{h},a))^{2})$
and plays the arg max action
$a_{h}\mathrel{\mathop{\ordinarycolon}}=\arg\max_{a}\tilde{Q}_{h}(s_{h},a).$
The algorithm then observes a reward $r_{h}$ and the next state $s_{h+1}$, computes a target $z$, and updates the posterior mean estimate using the Q-learning update rule. A natural setting of the target would be $r_{h}+\max_{a^{\prime}}\tilde{Q}_{h+1}(s_{h+1},a^{\prime})$, which we refer to as the “vanilla version". However, due to unresolvable difficulties in regret analysis discussed later in Section 4, the PSQL algorithm computes the target in a slightly optimistic manner as we describe next. Our experiments (Appendix A) show that although this modification does impact performance, PSQL still remains significantly superior to its UCB counterparts.
1:Input: Parameters: $\delta\in(0,1)$. Set $J\mathrel{\mathop{\ordinarycolon}}=J(\delta)$ as in (11).
2:Initialize: $\widehat{Q}_{H+1}(s,a)=\widehat{V}_{H+1}(s)=0$, $\widehat{Q}_{h}(s,a)=\widehat{V}_{h}(s)=H,N_{h}(s,a)=0\,\forall s,a,h$.
3:for episodes $k=1,2,\ldots$ do
4:  Observe $s_{1}$.
5:  for step $h=1,2,\ldots,H$ do
6:   Sample $\tilde{Q}_{h}(s_{h},a)\sim\mathcal{N}(\widehat{Q}_{h}(s_{h},a),\sigma(N_{h}(s_{h},a))^{2})$, for all $a\in\mathcal{A}$.
7:   Play $a_{h}\mathrel{\mathop{\ordinarycolon}}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\tilde{Q}_{h}(s_{h},a)$.
8:   Observe $r_{h}$ and $s_{h+1}$.
9:   $z\leftarrow\texttt{ConstructTarget}(r_{h},s_{h+1},\widehat{Q}_{h+1},N_{h+1})$.
10:   $n\mathrel{\mathop{\ordinarycolon}}=N_{h}(s_{h},a_{h})\leftarrow N_{h}(s_{h},a_{h})+1$, $\alpha_{n}\mathrel{\mathop{\ordinarycolon}}=\frac{H+1}{H+n}$.
11:   $\widehat{Q}_{h}(s_{h},a_{h})\leftarrow(1-\alpha_{n})\widehat{Q}_{h}(s_{h},a_{h})+\alpha_{n}z$.
12:  end for
13:end for
Algorithm 1 Q-learning with Posterior Sampling (PSQL)
1: Set $\widehat{a}=\operatorname*{arg\,max}_{a}\widehat{Q}(s^{\prime},a)+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma(N(s^{\prime},a))}$.
2: /* Take maximum of the $J$ samples from the posterior of target $V_{h+1}$ */
3:Sample $\tilde{V}^{j}\sim\mathcal{N}\mathinner{\left(\widehat{Q}(s^{\prime},\widehat{a}),\sigma(N(s^{\prime},\widehat{a}))^{2}\right)}$, for $j\in[J(\delta)]$, $a\in\mathcal{A}$ .
[…]
5:Return $z\mathrel{\mathop{\ordinarycolon}}=r+\overline{V}(s^{\prime})$.
Algorithm 2 ConstructTarget($r,s^{\prime},\widehat{Q},N)$
[…]
$\widehat{a}\mathrel{\mathop{\ordinarycolon}}=\arg\max_{a}\widehat{Q}_{h+1}(s^{\prime},a)+\sigma(N_{h+1}(s^{\prime},a))^{2},\text{ and}$
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\overline{V}_{h+1}(s^{\prime})=\max_{j\in J}\max_{a\in\mathcal{A}}\tilde{q}^{j}(a)$}
$\mbox{ with }\textstyle J=J(\delta)\coloneqq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{\log(SAT/\delta)}{\log(4/(4-p_{1}))}},\quad p_{1}=\Phi(-1)-\frac{\delta}{H}-\delta.$ (11)
Then,
$\overline{V}_{h+1}(s^{\prime})\mathrel{\mathop{\ordinarycolon}}=\max_{j\in[J]}\max_{a\in\mathcal{A}}\tilde{q}^{j}(a),\mbox{ and }z=r+\overline{V}(s^{\prime}).$
Observe that the above procedure computes a $\overline{V}_{h+1}(s^{\prime})$ that is more optimistic than single sample max ($\max_{a^{\prime}}\tilde{Q}_{h}(s^{\prime},a^{\prime})$). However, the optimism is limited only to the target computation not to the main decision-making in Line 6, marking an important departure from UCB based optimism (e.g., Jin et al. (2018)). The importance and rationale of this target computation procedure will be further explained in regret analysis (Section 4).
Finally, the algorithm uses the computed target $z$ to update the posterior mean via the Q-learning update rule, with $\alpha_{n}=\frac{H+1}{H+n}$:
$\widehat{Q}_{h}(s,a)\leftarrow(1-\alpha_{n})\widehat{Q}_{h}(s,a)+\alpha z.$
4 Regret Analysis
We prove the following regret bound for PSQL.
Theorem 2.
The cumulative regret of PSQL (Algorithm 1,2) in $K$ episodes satisfies
$\textstyle\text{Reg}(K)\mathrel{\mathop{\ordinarycolon}}=(\sum_{k=1}^{K}V^{*}_{1}(s_{k,1})-V_{1}^{\pi_{k}}(s_{k,1}))\leq O\mathinner{\left(H^{2}\sqrt{SAT}\chi\right)},$
with probability at least $1-\delta$, where $\chi=\log(JSAT/\delta)$ and $T=KH$.
4.1 Challenges and techniques.
Q-learning with posterior sampling poses several unique challenges for theoretical analysis.
Challenges under vanilla posterior sampling.
A key difficulty in analyzing Bayesian posterior sampling algorithms is the absence of high probability optimism (i.e., the property that the estimates upper bound the true parameters). To appreciate the importance of optimism, observe that the regret of an algorithm over $K$ episodes can be decomposed into two buckets:
$\textstyle\sum_{k=1}^{K}\mathinner{\left(V^{*}_{1}(s_{k,1})-V^{\pi_{k}}_{1}(s_{k,1})\right)}\leq\underbrace{\textstyle\sum_{k=1}^{K}(V^{*}_{1}(s_{k,1})-\tilde{Q}_{k,1}(s_{k,1},a_{k,1}))}_{\text{Optimism error}}+\underbrace{\textstyle\sum_{k=1}^{K}(\tilde{Q}_{k,1}(s_{k,1},a_{k,1})-V^{\pi_{k}}_{1}(s_{k,1}))}_{\text{Estimation error}}.$ (12)
In optimistic algorithms like UCBQL Jin et al. (2018), there is no optimism error since the UCB estimate is a high-confidence upper bound on the optimal value function. For posterior sampling, in the $N$-armed bandit setting, Agrawal and Goyal (2012, 2017) made a breakthrough by showing that optimism holds with constant probability in exploration phase, and is sufficient to bound the regret. However, the recursive nature of Q-learning update makes achieving even a constant probability of optimism non-trivial in Q-learning. To see this intuitively, suppose that if we have a constant probability $p$ of optimism of posteriors on value functions in state $H$. Then, due to the recursive nature of $Q$-value update, the optimism of stage H1H-1 value requires optimism of stage $H$ value; leading to $p^{2}$ probability of optimism in stage H1H-1. Continuing this way, we get an exponentially small probability of optimism for stage $1$.
Challenges under optimistic posterior sampling.
To get around the issues with constant probability optimism, many posterior sampling algorithms (e.g., model-based PSRL Agrawal and Jia (2022), MNL-bandit Agrawal et al. (2017),Tiapkin et al. (2022) etc.) taking max over multiple (say $J$) samples from the posterior in order to get high probability optimism. This modification is sometimes referred to as optimistic posterior sampling. We follow a similar modification, however, only for the target construction. We believe that, due to bootstrapping nature of Q-learning (or TD-learning methods in general) taking multiple samples for both decision-making and the target construction would lead to exponential accumulation of errors, even for $J$ as small as $2$. Below, we provide a rough intuitive argument.
We expect the bias of $\tilde{Q}$ (sample from the posterior distribution) to closely track $Q^{*}$. Now, suppose $J$ samples are used in decision-making (i.e., we take multiple samples from the posterior distributions at Line 6 of Algorithm 1). Consider step $h=H$, we would incur additional regret when the bias of decision-making sample, here, $\max_{j}\tilde{Q}^{j}$ exceeds $Q^{*}$ (optimism error). Using standard techniques, this error at step $H$ has an error bound of $\sigma\sqrt{J\log(1/\delta)}$ with probability $1-\delta$ where $\sigma^{2}$ is the variance of the distribution of $\tilde{Q}$. This error subsequently propagates multiplicatively via bootstrapping of Bellman equation.
For the step H1H-1, the optimism error contribution will be $\sigma\sqrt{J^{2}\log(1/\delta)}$ with probability $1-\delta$. Continuing this argument, we get that at step $1$, the cumulative optimism error will be of the order $\sigma\sqrt{J^{H}\log(1/\delta)}$, i.e., exponential in $H$ for any $J\geq 2$. To summarize, the usual trick of obtaining high probability optimism by taking multiple samples from the posterior doesn’t work for Q-learning, at least not without further novel ideas.
Our techniques.
To handle the above-mentioned challenges, our algorithm uses a combination of vanilla (single-sample) and optimistic (multiple-sample) posterior sampling for action selection and target computation, respectively. See Algorithm 1, 2 for details.
A key observation utilized here is that to break the recursive multiplicative decay in constant probability of optimism, we just need to ensure high probability optimism of the next-stage value function estimate used in the target computation. So, perhaps a solution would be to use optimistic posterior sampling in target computation only. That is, in the target computation, compute the next stage value function estimate as argmax over multiple samples from the posterior on Q-values, whereas the action taken is still decided using single sample posterior sampling. Indeed, this idea does ensure constant probability optimism while also resolving the issue of accumulating multiplicative ($J^{H}$) factors in variance and error bounds. However, now we run into the problem that the played action (single sample argmax of Q-values) may not match the action used in the target computation (multiple sample argmax). This poses difficulties in controlling the target estimation error, since the action used in the target computation may not have enough plays, and therefore, its variance may not decrease fast enough.
The final piece of the puzzle in our algorithm design and analysis is to obtain regret decomposition with only constant probability optimism for the posterior samples used for taking samples and the argmax action $\widehat{a}$ of the posterior mean (with a standard deviation offset) in our target computation. (For details, see Algorithm 2.) In our regret analysis, we show that this action $\widehat{a}$ either matches the played action with constant probability, or has a smaller variance (Lemma 5). In the next subsection, we provide a proof sketch that formally puts together all these ideas to obtain our main regret bound.
4.2 Proof sketch
We provide a proof sketch for Theorem 2. All the missing details from this section are in Appendix C. Here, we use $\tilde{Q}_{k,h},\widehat{Q}_{k,h},\overline{V}_{k,h},N_{k,h}$, to denote the values of $\tilde{Q}_{h},\widehat{Q}_{h},\overline{V}_{h},N_{h}$, respectively at the beginning of episode $k$ of Algorithm 1, 2. And, as before, $s_{k,h},a_{k,h}$ denote the state and action visited at episode $k$, step $h$.
Following the regret decomposition in (12), we bound the regret by bounding optimism error and estimation error. We first prove a constant probability optimism, so that the optimism error is 0 with a constant probability. Then, we build upon a technique introduced in Zanette et al. (2020); Agrawal et al. (2021) to convert this to a high probability optimism error bound. Obtaining such an optimistic error bound for stage $h$ requires an estimation error bound for stage $h+1$, which in turn uses optimism error bound for stage $h+1$. Therefore, the detailed proof in the appendix bounds estimation and optimism error inductively as one grand lemma (Lemma 8). In this proof sketch, we separate these bounds into abridged lemmas for a more intuitive exposition.
(a) Bounding optimism error.
We show optimism of the posterior sample $\tilde{Q}_{k,h}$ holds with constant probability in the following lemma.
Lemma 1 (Abridged).
[…]
$\tilde{Q}_{k,h}(s,a)\geq Q^{*}_{h}(s,a).$ (13)
The proof of the above lemma inductively uses the optimism and estimation error bounds available for the next stage ($h+1$) to bound the estimation error in the posterior mean $\widehat{Q}_{h}(s,a)$. Then, anti-concentration (i.e., lower tail bounds) of the Gaussian posterior distribution provides the desired constant probability optimism.
Next, building upon the technique from Zanette et al. (2020); Agrawal et al. (2021), we derive the following general lemma (proof in Appendix E), which we use to derive high-confidence bounds from constant probability bounds.
Lemma 2 (High confidence from constant probability).
For some fixed scalars $X^{*},\underline{X}$, and $p,\delta\in(0,1)$, suppose that $\tilde{X}\sim\mathcal{D}$ satisfies $\tilde{X}\geq X^{*}$ with probability at least $p$, $\tilde{X}\geq\underline{X}$ with probability at least $1-\delta$, and $\mathbb{E}[\tilde{X}]\geq\underline{X}$. Then, with probability at least $1-2\delta$,
$\textstyle\tilde{X}\geq X^{*}-\frac{1}{p}\mathinner{\left(\mathbb{E}_{\mathcal{D}}[\tilde{X}]-\underline{X}\right)}.$ (14)
We apply the above lemma with Lemma 1 to obtain a high probability optimism error bound.
Lemma 3 (Optimism error bound).
In Algorithm 1, with probability $1-2\delta$, the following holds for any $k\in[K]$ and $h\in[H]$,
$\textstyle\tilde{Q}_{k,h}(s_{k,h},a_{k,h})$ $\geq$ $\textstyle V^{*}_{h}(s_{k,h})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{2}{p_{1}}\sqrt{2\sigma(N_{k,h}(s_{k,h},a_{k,h}))^{2}\log(AKH/\delta)}},$
where $p_{1}$ is as defined in Lemma 1.
(b) Bounding estimation error.
In Q-learning, an estimate of the next stage value function (here, $\overline{V}_{h+1}$) is used to compute the target in order to update the $Q$-value for the current stage (here, the posterior mean $\widehat{Q}_{h}$). As a result, the error in the posterior mean for stage $h$ depends on the error in the value function estimates for $h+1$. This is essentially captured by the following lemma.
Lemma 4 (Posterior mean estimation error).
With probability at least $1-\delta$, for all $k,h,s,a\in[K]\times[H]\times\mathcal{S}\times\mathcal{A}$,
$\textstyle\widehat{Q}_{k,h}(s,a)-Q^{*}_{h}(s,a)\leq\sqrt{\sigma(N_{k,h}(s,a))^{2}\eta}+\alpha_{n}^{0}H+\sum_{i=1}^{n}\alpha_{n}^{i}\mathinner{\left(\overline{V}_{k_{i},h+1}(s_{k_{i},h+1})-V^{*}_{h+1}(s_{k_{i},h+1})\right)},$
where $n=N_{k,h}(s,a)$, and $\eta=\log(SAKH/\delta)$. And, $\alpha_{n}^{i}=\alpha_{i}\Pi_{j=i+1}^{n}(1-\alpha_{j}),\,i>0$, with $\alpha_{n}^{0}=\Pi_{j=1}^{n}(1-\alpha_{j})$.
Conceivably, we should be able to apply the above lemma inductively to obtain an estimation error bound (Lemma 6). However, a challenge here is that $\overline{V}_{k,h+1}(s_{k,h+1})$ is obtained by sampling from the posterior of $\tilde{Q}(s_{k,h+1},\cdot)$ at action $\widehat{a}$ ($\coloneqq\operatorname*{arg\,max}_{a}\widehat{Q}_{k,h}(s_{k,h},a)$) and not $a_{k,h+1}$ ($\coloneqq\operatorname*{arg\,max}_{a}\tilde{Q}_{k,h}(s_{k,h},a)$). So recursively applying this lemma would give bounds in terms of variance at $\widehat{a}$ and not $a_{k,h+1}$. The lemma below gets around this difficulty by showing that the variance at $\widehat{a}$ can be bounded in terms of the variance at $a_{k,h+1}$.
Lemma 5 (Bounding action mismatch).
For a given $k,h,s_{k,h}$, and let $\widehat{a}\coloneqq\operatorname*{arg\,max}_{a}\widehat{Q}_{k,h}(s_{k,h},a)+\sqrt{\sigma(N_{k,h}(s_{k,h},a))^{2}}$,
$\textstyle\sigma(N_{k,h}(s_{k,h},\widehat{a}))^{2}<2\sigma(N_{k,h}(s_{k,h},a_{k,h}))^{2}\log(1/\delta),\quad\text{with probability at least $p_{2}$},$
where $p_{2}=\Phi(-2)-(A)\delta$, and $\delta$ as defined in Theorem 2.
Lemma 6 (Cumulative estimation error.).
With probability at least $1-\delta$, the following holds for all $h\in[H]$,
$\textstyle\sum_{k=1}^{K}\mathinner{\left(\tilde{Q}_{k,h}(s_{k,h},a_{k,h})-V^{\pi_{k}}_{h}(s_{k,h})\right)}$ $\leq$ $O\mathinner{\left(H^{2}\sqrt{SAT}\log(JSAT/\delta)\right)}.$
(c) Putting it all together.
To obtain the regret bound in Theorem 2, we simply sum up the optimism error bound in Lemma 3 for $K$ episodes and add it to the cumulative estimation error bound provided in Lemma 6.
5 Conclusion
We presented a posterior sampling-based approach for incorporating exploration in Q-learning. Our PSQL algorithm is derived from an insightful Bayesian inference framework and shows promising empirical performance in preliminary experiments. (Detailed experimental setup and empirical results on additional environments are provided in Appendix A.) We proved a $\tilde{O}(H^{2}\sqrt{SAT})$ regret bound in a tabular episodic RL setting that closely matches the known lower bound. Future directions include a theoretical analysis of the vanilla version of PSQL that uses a single sample from next stage posterior in the target computation. The vanilla version outperforms Algorithm 1 empirically but is significantly harder to analyze. Another avenue is tightening the $H$ dependence in the regret bound; Appendix F outlines a sketch for improving it by $\sqrt{H}$ although at the expense of making the algorithm more complex. Further refinements are potentially achievable using techniques from Li et al. (2021a); Zhang et al. (2020).
Appendix A Experiments
In section 4, we proved that our Q-learning with posterior sampling algorithm  PSQL enjoys regret bounds comparable to its UCB-counterparts, e.g., Jin et al. [2018]. In this section, we present empirical results that validate our theory and compare the empirical performance of the posterior sampling approach against several benchmark UCB-based and randomized algorithms for reinforcement learning.
For the empirical studies, we use the vanilla version of posterior sampling, which we denote as PSQL* (see Figure 1). In this vanilla version, the target computation at a step $h$ is the default $z=r_{h}+\max_{a^{\prime}}\tilde{Q}_{h+1}(s_{h+1},a^{\prime})$. As discussed in Section 3.2, PSQL modified this target computation to make it slightly optimistic, to deal with the challenges in theoretical analysis. Later, we also compare the empirical performance of PSQL and PSQL*. While the modified target computation does slightly deteriorate the performance of PSQL, in our experiments, it still performs significantly better compared to the benchmark UCB-based approach Jin et al. [2018].
Specifically, we compare the posterior sampling approach to the following three alorithms.
  • UCBQL Jin et al. [2018] (Hoeffding version): the seminal work which gave the first UCB based Q-learning regret analysis.
  • RLSVI Russo [2019]: a popular randomized algorithm that implicitly maintains posterior distributions on Value functions.
  • Staged-RandQL Tiapkin et al. [2023]: a recently proposed randomized Q-learning based algorithm that uses randomized learning rates to motivate exploration.
Environment description.
We report the empirical performance of RL algorithms on two tabular environments described below. In each environment, we report each algorithm’s average performance over 10 randomly sampled instances.
  • (One-dimensional “chain” MDP:) An instance of this MDP is defined by two parameters $p\in[0.7,0.95]$ and $S\in\{7,8,9,\ldots,14\}$. In a random instance, p,&Sp,\&\,S are chosen randomly from the given ranges. The resultant MDP environment is a chain in which the agent starts at state 0 (the far-left state), and state $S$ (the right-most state) is the goal state. At any given step $h$ in an episode, the agent can take “left” or “right” action. The transitions are to the state in the direction of the action taken with probability $p$, and in the opposite direction with probability 1p1-p.
  • (Two-dimensional “grid-world” MDP, similar to FrozenLake environment in the popular Gymnasium library:) A random instance of this MDP is defined by a $4\times 4$ grid with a random number of “hole” states placed at on the grid uniformly at random that the agent must avoid or else the episode ends without any reward. The agent starts at the upper-left corner, and the goal state is the bottom-right corner of the grid. There is at least one feasible path from the starting state to the goal state that avoids all hole states. At any given time step, the agent can take the "left", "right", "bottom" and "up" actions. After an action is taken, the agent has $1/3$ probability to transit to the direction of the action taken, and $1/3$ probability each to transit to the two perpendicular directions.
In both the above environments, the goal state carries the reward of (Hh)/H(H-h)/H, where $H$ is the duration of the episode and $h$ is the time index within the episode at which the goal state is reached. No other state has any reward. The duration of an episode is set at $H=32$ for all experiments.
Findings.
We observed that the performance of all the algorithms is sensitive to constants in the exploration bonuses or in the posterior variances. These constants were tuned such that the respective algorithms performed the best in the two environments. We made the following parameter choices for the algorithmic simulations for a fair comparison:
  • $\delta$ is fixed for all algorithms as $0.05$.
  • In UCBQL Jin et al. [2018], the Q-function estimates are initialized as the maximum value of any state in the environment ($\eqqcolon V_{\max}$). The exploration bonus for any $h,s,a$ with visit counts as $n$ is given by
    $\sqrt{c\frac{V^{2}_{\max}\log(SAT/\delta)}{n}},$
    with $c=0.01$
  • In PSQL and PSQL*, the Q-function posterior means are initialized as $V_{\max}$ (same as UCBQL) and the standard deviation of the posterior for any $h,s,a$ with $n$ visits is given by
    $\sqrt{c\frac{V^{2}_{\max}}{\max\{1,n\}}},$
    with $c=0.02$.
  • In RLSVI Russo [2019], the per-reward perturbation is a mean zero Gaussian with standard deviation for any $h,s,a$ with $n$ visits is given by,
    $\sqrt{c\frac{V^{2}_{\max}\log(SAT/\delta)}{n+1}},$
    with $c=0.005$.
  • In Staged-RandQL Tiapkin et al. [2023], for the initialization of the Q-function estimates, we use a tighter upper bound of (Hh)/H(H-h)/H at step $h$ available in our environment, instead of the default HhH-h suggested in their paper. We use $n_{0}=1/S$ and $r_{0}=1$ as in their paper.
Our results are summarized in Figure 2 and 3. The error bars represent one standard deviation interval around the mean cumulative regret of an algorithm over 10 runs on randomly generated instances of the environment.
(a) One-dimensional “chain” MDP
(b) Two-dimensional “grid” MDP
Figure 2: Regret comparison: x-axes denotes episode index, y-axes denotes cumulative regret
We observe that the randomized/posterior sampling algorithms PSQL*, RLSVI, and Staged-RandQL, have lower regret than their UCB counterpart: UCBQL. Also, PSQL* has significantly lower regret than the other two randomized algorithms.
A direct practical implication is that, PSQL* enjoys a shorter learning time (number of episodes after which the cumulative regret is below the specified threshold [Osband et al., 2019]). Further, the variance across different runs is also the lowest of all, suggesting PSQL* enjoys higher robustness.
In Figure 3, we compare the performance of PSQL*, the single sample vanilla version of posterior sampling, with the PSQL algorithm for which we provided regret bounds. As we explained in Section 4.1 (Challenges and Techniques), in order to achieve optimism in the target, PSQL computed the next state value by taking the max over multiple samples from the posterior of empirical mean maximizer action $\widehat{a}=\arg\max_{a}\widehat{Q}_{h}(s_{h},a)$. This introduces some extra exploration, and as a result, we observe that PSQL* displays a more efficient exploration-exploitation tradeoff, PSQL still performs significantly better than the UCB approach. These observations motivate an investigation into the theoretical analysis of the vanilla version, which we believe will require significantly new techniques.
(a) One-dimensional “chain” MDP
(b) Two-dimensional “grid” MDP
Figure 3: Regret comparison: x-axes denotes episode index, y-axes denotes cumulative regret
Appendix B Bayesian inference based interpretation for Q-learning
In this section, we describe the mathematical steps for calculating the updated posterior distribution from (8).
First, in Proposition 1, we derive the well-known result that solving the optimization problem in (6) gives the posterior distribution as expected by Bayes rule. Let $\theta\in\Theta$ be the Bayesian parameter that we are inferring with $\Delta_{\Theta}$ be the space of distributions on $\Theta$. Let $p(\theta)$, $\ell(\theta,\cdot)$, and $q(\theta)$ be the the current prior distribution on $\theta$, the negative log likelihood function and the posterior distribution to be calculated.
Proposition 1 (Also in Khan and Rue [2023], Knoblauch et al. [2022]).
Let $KL(q(\theta)||p(\theta))=\int_{\theta}q(\theta)\log(\frac{q(\theta)}{p(\theta)})$. Given log likellihood function $\ell(\theta,z)$, and prior $p(\theta)$, the distribution $q$ that maximizes ELBO objective,
$\max_{q\in\Delta_{\Theta}}\mathbb{E}_{\theta\sim q}[\ell(\theta,z)]-KL(q(\theta)||p(\theta))$ (15)
is given by the Bayes rule
$q^{\text{Bayes}}(\theta)\propto p(\theta)\cdot\exp(\ell(\theta,z)).$ (16)
Proof.
Note that the ELBO objective function is equivalent to
$-\int_{\theta}\log(\exp(-\ell(\theta,z)q(\theta)-\int_{\theta}\log\mathinner{\left(\frac{q(\theta)}{p(\theta)}\right)}q(\theta)$
$=$ $-\int_{\theta}\log\mathinner{\left(\frac{q(\theta)}{p(\theta)\exp(\ell(\theta,z))}\right)}q(\theta),$
which is maximized when $q(\theta)=q^{\text{Bayes}}(\theta)$. ∎
Now, we study the calculation of the posterior distribution of $Q_{h}(s,a)$ after observing $n+1$ visits of $(h,s,a)$ in Lemma 7.
Lemma 7.
Consider the following maximization problem (regularized ELBO) over the space $\Delta_{\Theta}$ of distributions over a parameter $\theta$.
$\max_{q\in\Delta_{\Theta}}\mathbb{E}_{\theta\sim q}[\ell(\theta,z)]-KL(q(\theta)||p(\theta))+\lambda_{n}{\cal H}(q(\theta)),$ (17)
Then, if $p(\cdot)$ is given by the pdf of the Gaussian distribution ${\cal N}(\widehat{\mu}_{n-1},\frac{\sigma^{2}}{n-1})$, and $\ell(\theta,z)=\log(\phi_{\theta}(z))$ where $\phi_{\theta}(z)=\Pr(z|\theta)$ is the pdf of the Gaussian distribution $\mathcal{N}(\theta,\frac{\sigma^{2}}{H+1})$, and $\lambda_{n}=\frac{H}{n}$; then the optimal solution $q(\cdot)$ to (17) is given by the Gaussian distribution ${\cal N}(\widehat{\mu}_{n},\frac{\sigma^{2}}{n})$, where
$\widehat{\mu}_{n}=(1-\alpha_{n})\widehat{\mu}_{n-1}+\alpha_{n}z,\text{ with }\alpha_{n}=\frac{H+1}{H+n}.$
Proof.
Denote the objective value at a given distribution $q$ as $rELBO(q)$. Then,
$rELBO(q)$ $=$ $\int_{\theta}\log(\exp(\ell(\theta,z)q(\theta)-\int_{\theta}\log\mathinner{\left(\frac{q(\theta)}{p(\theta)}\right)}q(\theta)-\lambda_{n}\int_{\theta}\log(q(\theta))q(\theta)$
$=$ $-\int_{\theta}\log\mathinner{\left(\frac{q(\theta)^{1+\lambda_{n}}}{p(\theta)\exp(\ell(\theta,z))}\right)}q(\theta),$
which is maximized at distribution $q$ with $q(\theta)\propto(p(\theta)\exp(\ell(\theta,z)))^{1/(\lambda_{n}+1)}$ Then,
$q(\theta)$ $\propto$ $\exp\mathinner{\left[\frac{1}{\lambda_{n}+1}\mathinner{\left(-\frac{(n-1)\mathinner{\left(\theta-\widehat{\mu}_{n-1}\right)}^{2}}{2\sigma^{2}}-\frac{(H+1)(z-\theta)^{2}}{2\sigma^{2}}\right)}\right]}$
$\propto$ $\exp\mathinner{\left[-\frac{n}{H+n}\mathinner{\left(\frac{\theta^{2}(H+n)-2\theta((n-1)\widehat{\mu}_{n-1}+(H+1)z)}{2\sigma^{2}}\right)}\right]}$
$\propto$ $\exp\mathinner{\left[-n\mathinner{\left(\frac{\theta^{2}-2\theta\widehat{\mu}_{n}}{2\sigma^{2}}\right)}\right]}$
$\propto$ $\exp\mathinner{\left[-n\mathinner{\left(\frac{(\theta-\widehat{\mu}_{n})^{2}}{2\sigma^{2}}\right)}\right]}$
where $\widehat{\mu}_{n}=\frac{n-1}{H+n}\widehat{\mu}_{n-1}+\frac{H+1}{H+n}z=(1-\alpha_{n})\widehat{\mu}_{n-1}+\alpha_{n}z$. ∎
Appendix C Missing Proofs from Section 4
C.1 Optimism
Lemma 8 (Unabridged version).
The samples from the posterior distributions and the mean of the posterior distributions as defined in Algorithm 12 satisfy the following properties: for any episode $k\in[K]$ and index $h\in[H]$,
  1. (a)
    (Posterior distribution mean) For any given $s,a$, with probability at least $1-\frac{\delta}{H}-\delta$,
    $\widehat{Q}_{k,h}(s,a)\geq Q^{*}_{h}(s,a)-\sqrt{\sigma(N_{k,h}(s,a))^{2}}.$ (18)
  2. (b)
    (Posterior distribution sample) For any given $s,a$, with probability at least $p_{1}$ ($p_{1}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Phi(-1)-\frac{\delta}{H}-\delta}$),
    $\tilde{Q}_{k,h}(s,a)\geq Q^{*}_{h}(s,a).$ (19)
  3. (c)
    (In Algorithm 2) With probability at least $1-\frac{\delta}{KH}$,
    $\overline{V}_{k,h}(s_{k,h})\geq V^{*}_{h}(s_{k,h}).$ (20)
Here $\delta$ is a parameter of the algorithm used to define the number of samples $J$ used to compute the target $\overline{V}$.
Proof.
We prove the lemma statement via induction over $k,h$.
Base case: $k=1,h\in[H]$.
Note that $n_{1,h}(s,a)=0$ for all $s,a,h\in\mathcal{S}\times\mathcal{A}\times[H]$. Therefore, $\widehat{Q}_{1,h}(s,a)=H$ for all $s,a,h$ ((18) is trivially true). As $Q^{*}_{h}(s,a)\leq H$ for all $s,a,h$, therefore $\tilde{Q}_{1,H}(s,a)\geq Q^{*}_{H}(s,a),$ with probability at least $1/2$ ($p_{1}$), i.e. (19) is true. By the choice of $J$ and Lemma 13, (20) also follows.
Induction hypothesis:
Given $k>1,1\leq h\leq H$, assume that the statements (a),(b), and (c) are true for $1\leq k^{\prime}\leq k-1,h^{\prime}\in[H]$, and for $k^{\prime}=k,h+1\leq h^{\prime}\leq H$ .
Induction step:
For $k,h$, we show (18) holds with probability $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1-\frac{\delta}{H}-\delta$}, (19) holds with probability at least $p_{1}=\Phi(-1)-\frac{\delta}{H}-\delta$, and finally (20) holds with probability $1-\delta/KH$.
In case $n_{k,h}(s,a)=0$, then $\widehat{Q}_{k,h}(s,a)=H$ and $b_{k,h}(s,a)>0$ and therefore by the same reasoning as in the base case, the induction statement holds. For the rest of the proof we consider $n_{k,h}(s,a)>0$.
Let $n=N_{k,h}(s,a)$, then Algorithm 1 implies,
$\widehat{Q}_{k,h}(s,a)=\alpha_{n}^{0}H+\sum_{i=1}^{n}\alpha_{n}^{i}\mathinner{\left(r_{k_{i},h}+\overline{V}_{k_{i},h+1}(s_{k_{i},h+1})\right)},$ (21)
To prove the induction step for (18), consider the following using (21) and Bellman optimality equation.
$\widehat{Q}_{k,h}(s,a)-Q^{*}_{h}(s,a)$ $=$ $\sum_{i=1}^{n}\alpha_{n}^{i}\mathinner{\left(r_{k_{i},h}-r_{h}(s,a)+\overline{V}_{k_{i},h+1}(s_{k_{i},h+1})-P_{h,s,a}V^{*}_{h}\right)}$ (22)
$=$ $\sum_{i=1}^{n}\alpha_{n}^{i}\mathinner{\left(r_{k_{i},h}-r_{h}(s,a)+V^{*}_{h+1}(s_{k_{i},h+1})-P_{h,s,a}V^{*}_{h}\right)}$
$+\sum_{i=1}^{n}\alpha_{n}^{i}\mathinner{\left(\overline{V}_{k_{i},h+1}(s_{k_{i},h+1})-V^{*}_{h+1}(s_{k_{i},h+1})\right)}$
(using  (20) from induction hypothesis for $h+1\leq H$, taking union bound over $k$)
(with probability $1-\delta/H$)
(note that this is trivially true when $h+1=H+1$ since $\overline{V}_{k,H+1}=V^{*}_{h+1}=0$)
$\geq$ $\sum_{i=1}^{n}\alpha_{n}^{i}\mathinner{\left(r_{k_{i},h}-r_{h}(s,a)+V^{*}_{h+1}(s_{k_{i},h+1})-P_{h,s,a}V^{*}_{h}\right)}$
(using Corollary 4 with probability $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1-\delta$})
$\geq$ $-4\sqrt{\frac{H^{3}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log(1/\delta)}}{n_{k,h}(s,a)+1}}$
$=$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\sqrt{\sigma(N_{k,h}(s,a))^{2}}},$
Therefore, with a union bound we have with probability $1-\frac{\delta}{H}-\delta$,
$\widehat{Q}_{k,h}(s,a)$ $\geq$ $Q^{*}_{h}(s,a)-\sqrt{\sigma(N_{k,h}(s,a))^{2}}$ (23)
From the definition of cumulative density of Gaussian distribution we get,
$\text{Pr}\mathinner{\left(\tilde{Q}_{k,h}(s,a)\geq\widehat{Q}_{k,h}(s,a)+\sqrt{\sigma(N_{k,h}(s,a))^{2}}\right)}\geq\Phi(-1).$
Using (23), we have
$\text{Pr}\mathinner{\left(\tilde{Q}_{k,h}(s,a)\geq Q^{*}_{h}(s,a)\right)}$ $\geq$ $\Phi(-1)-\frac{\delta}{H}-\delta\eqqcolon p_{1}.$
Now, we show $\tilde{Q}_{k,h}(s,\widehat{a})\geq Q^{*}_{h}(s,a^{*})=V^{*}_{h}(s)$ with probability at least $\Phi(-1)-\delta-\delta/H$, where
$\widehat{a}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widehat{Q}_{k,h}(s,a)+\sigma(N(s,a)),\text{ and }a^{*}=\operatorname*{arg\,max}_{a\in\mathcal{A}}Q^{*}_{h}(s,a).$
By the definition of $\widehat{a}$ and properties of Gaussian distribution:
[…]
Therefore, $\tilde{Q}_{k,h}(s,\widehat{a}\geq Q^{*}_{k,h}(s,a^{*})$ with probability $\Phi(-1)-\delta-\delta/H=p_{1}$.
Setting $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}J\geq\frac{\log(KH/\delta)}{\log(1/(1-p_{1}))}$} in Algorithm 1, we use Lemma 13 to prove (20) holds with probability at least $1-\delta/KH$.
C.2 Action mismatch bound
See 5
Proof.
Consider the following partition of $A$ actions:
$\overline{\mathcal{A}}$ $\coloneqq$ $\mathinner{\left\{a\mathrel{\mathop{\ordinarycolon}}2\sigma(N(s,a))^{2}\log(1/\delta)>\sigma(N(s,\widehat{a}))^{2}\right\}},$
$\underline{\mathcal{A}}$ $\coloneqq$ $\mathinner{\left\{a\mathrel{\mathop{\ordinarycolon}}2\sigma(N(s,a))^{2}\log(1/\delta)\leq\sigma(N(s,\widehat{a}))^{2}\right\}}.$
Clearly, $\widehat{a}\in\overline{\mathcal{A}}$. We prove that with probability at least $\Phi(-2)-A\delta$, we have $\tilde{a}\in\overline{\mathcal{A}}$, so that $\sigma(N(s,\widehat{a}))^{2}<2\sigma(N(s,a))^{2}\log(1/\delta)$.
By definition of $\widehat{a}$, we have that for all $a$,
$\widehat{Q}(s,a)+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sqrt{\sigma(N(s,a))^{2}\log(JA/\delta)}}\leq\widehat{Q}(s,\widehat{a}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\sqrt{\sigma(N(s,\widehat{a}))^{2}\log(JA/\delta)}}.$
Also, by construction of $\underline{\cal A}$, we have that for $\forall a\in\underline{\mathcal{A}}$,
$\widehat{Q}(s,a)+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sqrt{\sigma(N(s,a))^{2}}}+\sqrt{2\sigma(N(s,a))^{2}\log(1/\delta)}\leq\widehat{Q}(s,\widehat{a}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\sqrt{\sigma(N(s,\widehat{a}))^{2}}}+\sqrt{\sigma(N(s,\widehat{a}))^{2}}$
From Gaussian tail bounds (see Corollary 3) we have for $\forall a\in\mathcal{A}$,
$\text{Pr}\mathinner{\left(\tilde{Q}(s,a)\leq\widehat{Q}(s,a)+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sqrt{\sigma(N(s,a))^{2}}}+\sqrt{2\sigma(N(s,a))^{2}\log(1/\delta)}\right)}\geq 1-A\delta$
so that $\forall a\in\underline{\mathcal{A}}$,
$\text{Pr}\mathinner{\left(\tilde{Q}(s,a)\leq\widehat{Q}(s,\widehat{a})+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2}\sqrt{\sigma(N(s,\widehat{a}))^{2}}\right)}\geq 1-A\delta$
[…]
$\tilde{Q}(s,\widehat{a})>\widehat{Q}(s,\widehat{a})+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2}\sqrt{\sigma(N(s,\widehat{a}))^{2}},$
Using a union bound on the last two events,, we get
$\tilde{Q}(s,\widehat{a})>\max_{a\in\underline{\mathcal{A}}}\tilde{Q}(s,a),\quad\text{with probability at least }\Phi(-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2})-A\delta.$ (24)
Since $\widehat{a}$ is in $\overline{\mathcal{A}}$, in the above scenario, the action $\tilde{a}$ that maximizes $\tilde{Q}(s,\cdot)$ must be in $\overline{\mathcal{A}}$. Therefore, $\sigma(N(s,\widehat{a}))^{2}<2\sigma(N(s,a))^{2}\log(1/\delta)$ with probability $\Phi(-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2})-A\delta$.
Proposition 2 (High probability action mismatch).
For a given $k,h,s_{k,h}$, let $\widehat{a}\coloneqq\operatorname*{arg\,max}_{a}\widehat{Q}_{k,h}(s_{k,h},a)$. Then,
$\sigma(N_{k,h}(s_{k,h},\widehat{a}))<\sigma(N_{k,h}(s_{k,h},a_{k,h}))\sqrt{2\log(1/\delta)}+\frac{1}{p_{2}}\mathbb{E}_{a_{k,h}}[\sigma(N_{k,h}(s_{k.h},a_{k,h}))\sqrt{2\log(1/\delta)}].$
where $p_{2}=\Phi(-2)-(A)\delta$, and $\delta$ as defined in Theorem 2. Here $\mathbb{E}_{a_{k,h}}[\cdot]$ denotes expectation over $a_{k,h}$ given $s_{k,h}$ and the history before round $k,h$.
Proof.
This follows by using previous lemma along with Lemma 2 with $\tilde{X}=\sigma(N_{k,h}(s_{k,h},a_{k,h}))\sqrt{2\log(1/\delta)}$, $)X^{*}=\sigma(N_{k,h}(s_{k,h},\widehat{a}$)), and $\underline{X}=0$. ∎
C.3 Target estimation error bound
The following lemma characterizes the target estimation error.
Lemma 9 (Target estimation error).
In Algorithm 1, with probability $1-2\delta$, the following holds for all $k\in[K]$ and $h\in[H]$,
$\overline{V}_{k,h}(s_{k,h})-\tilde{Q}_{k,h}(s_{k,h},a_{k,h})$
$\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}4\sigma(N_{k,h}(s_{k,h},a_{k,h}))\log(JKH/\delta)+\frac{4}{p_{2}}\mathbb{E}_{k,h}[\sigma(N_{k,h}(s_{k,h},a_{k,h}))]\log(JKH/\delta)}$
$\eqqcolon$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{1}{p_{2}}F(k,h,\delta),}$
where $p_{2}=\Phi(-2)-(A)\delta$ and $\delta$ as defined in Theorem 2, $J$ is defined in (11), and $\mathbb{E}_{a}\mathinner{\left[\cdot\right]}$ denotes expectation over the randomness in the action taken at $k,h$ conditioned on all history at the start of the $h_{\text{th}}$ step in the episode $k$ (i.e., only randomness is that in the sampling from the posterior distribution).
Proof.
[…]
$\overline{V}(s)$ $\leq$ $\widehat{Q}(s,\widehat{a})+\sqrt{2\sigma(N(s,\widehat{a}))^{2}\log(J/\delta)}$
$\leq$ $\tilde{Q}(s,\widehat{a})+\sqrt{2\sigma(N(s,\widehat{a}))^{2}\log(J/\delta)}+\sqrt{2\sigma(N(s,\widehat{a}))^{2}\log(1/\delta)}$
[…]
$\overline{V}(s)$ $\leq$ $\tilde{Q}(s,\widehat{a})+2\sqrt{2\sigma(N(s,\widehat{a}))^{2}\log(J/\delta)}$ (25)
$\leq$ $\tilde{Q}(s,\tilde{a})+2\sqrt{2\sigma(N(s,\widehat{a}))^{2}\log(J/\delta)}.$
To complete the proof, we use Proposition 2, and an union bound over all $k,h$ to have the following with probability at least $1-2\delta$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\overline{V}_{k,h}(s_{k,h})-\tilde{Q}_{k,h}(s_{k,h},a_{k,h})}$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\leq 4\sigma(N_{k,h}(s_{k,h},a_{k,h}))\sqrt{\log(JKH/\delta)\log(KH/\delta)}+\frac{4}{p_{2}}\mathbb{E}_{k,h}[\sigma(N_{k,h}(s_{k,h},a_{k,h}))]\sqrt{\log(JKH/\delta)\log(KH/\delta)}}.$
C.4 Optimism error bound
Corollary 1 (Optimism error bound).
In Algorithm 1, with probability $1-3\delta$, the following holds for any $k\in[K]$ and $h\in[H]$,
$V^{*}_{h}(s_{k,h})-\tilde{Q}_{k,h}(s_{k,h},a_{k,h})$
$\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}4\sigma(N_{k,h}(s_{k,h},a_{k,h}))\log(JKH/\delta)+\frac{4}{p_{2}}\mathbb{E}_{k,h}[\sigma(N_{k,h}(s_{k,h},a_{k,h}))]\log(JKH/\delta)}$
$\eqqcolon$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{1}{p_{2}}F(k,h,\delta),}$
where $p_{2}=\Phi(-2)-(A)\delta$ and $\delta$ as defined in Theorem 2, $J$ is defined in (11), and $\mathbb{E}_{a}\mathinner{\left[\cdot\right]}$ denotes expectation over the randomness in the action taken at $k,h$ conditioned on all history at the start of the $h_{\text{th}}$ step in the episode $k$ (i.e., only randomness is that in the sampling from the posterior distribution).
Proof.
From Lemma 9, we have with probability at least $1-2\delta$,
$\overline{V}_{k,h}(s_{k,h})-\tilde{Q}_{k,h}(s_{k,h},a_{k,h})$
$\leq 4\sigma(N_{k,h}(s_{k,h},a_{k,h}))\log(JKH/\delta)+\frac{4}{p_{2}}\mathbb{E}_{k,h}[\sigma(N_{k,h}(s_{k,h},a_{k,h}))]\log(JKH/\delta)$
$\eqqcolon$ $\frac{1}{p_{2}}F(k,h,\delta).$
Further, Lemma 8 (c) gives with probability at least $1-\delta$,
$V^{*}_{h}(s_{k,h})\leq\overline{V}_{k,h}(s_{k,h}).$
We complete the proof via a union bound.
Corollary 2.
With probability $1-\delta$, the following holds for all $h\in[H]$,
$\sum_{k=1}^{K}V^{*}_{h}(s_{k,h})-\tilde{Q}_{k,h}(s_{k,h},a_{k,h})$ $\leq$ $\mathrm{O}\mathinner{\left(\sqrt{H^{2}SAT}\chi\right)},$
where $\chi=\log(JSAT/\delta)$.
Proof.
From Corollary 1, we have for all $k,h$ simultaneously, with probability $1-3\delta$,
$\overline{V}_{k,h}(s_{k,h})-\tilde{Q}_{k,h}(s_{k,h},a_{k,h})$
$\leq 4\sigma(N_{k,h}(s_{k,h},a_{k,h}))\log(JKH/\delta)+\frac{4}{p_{2}}\mathbb{E}_{k,h}[\sigma(N_{k,h}(s_{k,h},a_{k,h}))]\log(JKH/\delta).$
By combining the definition of the variance in (10) with Corollary 6, the result follows easily.
C.5 Posterior mean estimation error bound
See 4
Proof.
First consider a fixed $k,h,s,a$. From (21) and Bellman optimality equation, we have (assume $n=N_{k,h}(s,a)\geq 1$),
$\widehat{Q}_{k,h}(s,a)-Q^{*}_{h}(s,a)$ $=$ $\sum_{i=1}^{n}\alpha_{n}^{i}\mathinner{\left(r_{k_{i},h}-r_{h}(s,a)+\overline{V}_{k_{i},h+1}(s_{k_{i},h+1})-P_{h,s,a}V^{*}_{h}\right)}$
$=$ $\sum_{i=1}^{n}\alpha_{n}^{i}\mathinner{\left(r_{k_{i},h}-r_{h}(s,a)+V^{*}_{h+1}(s_{k_{i},h+1})-P_{h,s,a}V^{*}_{h}\right)}$
$+\sum_{i=1}^{n}\alpha_{n}^{i}\mathinner{\left(\overline{V}_{k_{i},h+1}(s_{k_{i},h+1})-V^{*}_{h+1}(s_{k_{i},h+1})\right)}$
(Using Corollary 3 with probability $1-\delta$)
$\leq$ $\sqrt{\sigma(N_{k,h}(s,a))^{2}\log(1/\delta)}+\sum_{i=1}^{n}\alpha_{n}^{i}\mathinner{\left(\overline{V}_{k_{i},h+1}(s_{k_{i},h+1})-V^{*}_{h+1}(s_{k_{i},h+1})\right)}.$
When $N_{k,h}(s,a)=0$, then trivially $\widehat{Q}_{k,h}(s,a)-Q^{*}_{h}(s,a)\leq H=\alpha_{n}^{0}H$, and for $N_{k,h}(s,a)>0$, then $\alpha_{n}^{0}=0$. Combining these two cases and with a union bound over all $s,a,h,k$, we complete the proof. ∎
C.6 Cumulative estimation error bound
See 6
Proof.
For the purpose of writing this proof, define $\phi_{k,h}\coloneqq\tilde{Q}_{k,h}(s_{k,h},a_{k,h})-V^{*}_{h}(s_{k,h})$, $\delta_{k,h}\coloneqq\tilde{Q}_{k,h}(s_{k,h},a_{k,h})-V^{\pi_{k}}_{h}(s_{k,h})$, and $\beta_{k,h}\coloneqq\overline{V}_{k,h}(s_{k,h})-V^{*}_{h}(s_{k,h})$. Clearly $\delta_{k,h}\geq\phi_{k,h}$. Further, $v(n_{k,h})\leftarrow\sigma(N_{k,h}(s_{k,h},a_{k,h}))^{2}$.
Now consider,
$\tilde{Q}_{k,h}(s_{k,h},a_{k,h})-V^{\pi_{k}}_{h}(s_{k,h})$ $\leq$ $\tilde{V}_{k,h}(s_{k,h})-Q^{\pi_{k}}_{h}(s_{k,h},a_{k,h})$ (27)
$=$ $\tilde{Q}_{k,h}(s_{k,h},a_{k,h})-Q^{*}_{k,h}(s_{k,h},a_{k,h})+Q^{*}_{k,h}(s_{k,h},a_{k,h})-Q^{\pi_{k}}_{1}(s_{k,1},a_{k,1})$
(from Lemma 4 with probability $1-\delta$, with $n\leftarrow N_{k,h}(s_{k,h},a_{k,h})$)
$\leq$ $\alpha_{n}^{0}H+\sqrt{2v(n_{k,h})\eta}+\sum_{i=1}^{n}\alpha^{i}_{n}\beta_{k_{i},h+1}+P_{s_{k,h},a_{k,h}}\cdot(V^{*}_{h+1}-V^{\pi_{k}}_{h+1})$
$=$ $\alpha_{n}^{0}H+\sqrt{2v(n_{k,h})\eta}+\sum_{i=1}^{n}\alpha^{i}_{n}\beta_{k_{i},h+1}-\phi_{k,h+1}+\delta_{k,h+1}$
$+P_{s_{k,h},a_{k,h}}\cdot(V^{*}_{h+1}-V^{\pi_{k}}_{h+1})-(V^{*}_{h+1}(s_{k,h+1})-V^{\pi_{k}}_{h+1}(s_{k,h+1}))$
From Lemma 9 with probability $1-2\delta$
$\leq$ $\alpha_{n}^{0}H+\sqrt{2v(n_{k,h})\eta}+\frac{1}{p_{2}}\sum_{i=1}^{n}\alpha_{n}^{i}F(k_{i},h+1,\delta)+m_{k,h}$
$+\sum_{i=1}^{n}\alpha^{i}_{n}\phi_{k_{i},h+1}-\phi_{k,h+1}+\delta_{k,h+1},$
where
$m_{k,h}\coloneqq P_{s_{k,h},a_{k,h}}\cdot(V^{*}_{h+1}-V^{\pi_{k}}_{h+1})-(V^{*}_{h+1}(s_{k,h+1})-V^{\pi_{k}}_{h+1}(s_{k,h+1})).$
Now, we club all episodes together to have,
$\sum_{k=1}^{K}\delta_{k,h}$ $\leq$ $\sum_{k=1}^{K}\alpha_{n}^{0}H+\sum_{k=1}^{K}\sqrt{4v(n_{k,h})\eta}+\frac{1}{p_{2}}\sum_{k=1}^{K}\sum_{i=1}^{n}\alpha_{n}^{i}F({k_{i},h+1},\delta/KH))+\sum_{h=1}^{H}\sum_{k=1}^{K}m_{k,h}$
$+\sum_{k=1}^{K}\sum_{i=1}^{n}\alpha^{i}_{n}\phi_{k_{i},h+1}-\sum_{k=1}^{K}\phi_{k,h+1}+\sum_{k=1}^{K}\delta_{k,h+1}.$
From Lemma 16 (c), it follows ($a_{k,h}=\{\phi_{k,h},F(k,h,\delta/KH)\})$ $\sum_{k=1}^{K}\sum_{i=1}^{n}\alpha_{n}^{i}a_{k,h+1}\leq(1+1/H)\sum_{k=1}^{K}a_{k,h+1}$. Further from the initialization,
$\sum_{k=1}^{K}\alpha_{n_{k,h}}^{0}H\leq\sum_{k=1}^{K}\mathbb{I}\{n_{k,h}=0\}H\leq SAH=HSA.$
Therefore, we have,
$\sum_{k=1}^{K}\delta_{k,h}$ $\leq$ $\sum_{k=1}^{K}SAH+\sum_{k=1}^{K}\mathinner{\left(\sqrt{4v(n_{k,h})\eta}+m_{k,h}\right)}+(1+1/H)\frac{1}{p_{2}}\sum_{k=1}^{K}F({k,h+1},\delta/KH))+$
$+(1+1/H)\sum_{k=1}^{K}\phi_{k,h+1}-\sum_{k=1}^{K}\phi_{k,h+1}+\sum_{k=1}^{K}\delta_{k,h+1}.$
Unrolling the above $H$ times to have with a union bound over $k,h\in[K]\times[H]$ with probability $1-\delta$ ($\delta$ is scaled by $1/KH$ due to the union bound):
$\sum_{k=1}^{K}\tilde{Q}_{k,1}(s_{k,1},a_{k,1})-V^{\pi_{k}}_{1}(s_{k,1})$ $\leq$ $eSAH^{2}+e\sum_{h=1}^{H}\sum_{k=1}^{K}\mathinner{\left(\sqrt{4v(n_{k,h})\eta}+m_{k,h}\right)}$ (28)
$+\frac{e}{p_{2}}\sum_{h=1}^{H}\sum_{k=1}^{K}F({k,h+1},\delta/KH)),$
where we have used $\delta_{k,h}\geq\phi_{k,h}$ and $\delta_{k,H+1}=0$. Now, we analyze each term on the right hand side of (28) one by one. From Corollary 5,
$\sum_{h=1}^{H}\sum_{k=1}^{K}\sqrt{4v(n_{k,h})\eta}$ $\leq$ $\mathrm{O}\mathinner{\left(\sqrt{H^{4}SAT\eta}\right)}.$
From Corollary 6 with probability at least $1-\delta$,
$\frac{1}{p_{2}}\sum_{h=1}^{H}\sum_{k=1}^{K}\sum_{i=1}^{n}F(k_{i},h+1,\delta/KH)$ $\leq$ $\frac{1+H}{Hp_{2}}\sum_{h=1}^{H}\sum_{k=1}^{K}F(k,h+1,\delta/KH)$ (29)
$\leq$ $O(1)\cdot\sum_{h=1}^{H}\sum_{k=1}^{K}\sqrt{v(n_{k,h})\chi\log(KH/\delta)}$
$\leq$ $O\mathinner{\left(\sqrt{H^{4}SAT}\chi\right)},$ (30)
where $\chi=\log(JSAT/\delta)$.
Finally, from Lemma 10, we have with probability $1-\delta$
$\sum_{h=1}^{H}\sum_{k=1}^{K}m_{k,h}$ $\leq$ $\mathrm{O}\mathinner{\left(\sqrt{H^{4}T\log(KH/\delta)}\right)}.$
Combining the above, we complete the proof. ∎
See 2
Proof.
First we combine Corollary 2 and Lemma 6 to get with probability at least $1-\delta$ ($\chi=\log(JSAT/\delta)$),
$\sum_{k=1}^{K}V^{*}_{1}(s_{k,1})-V^{\pi_{k}}_{1}(s_{k,1})\leq O\mathinner{\left(\sqrt{H^{4}SAT}\chi\right)}.$
Observing that the rewards are bound, there $|\sum_{k=1}^{K}\sum_{h=1}^{H}R_{h}(s_{k,h},a_{k,h})-\sum_{k=1}^{K}V^{\pi_{k}}_{1}(s_{k,1})|\leq O(H\sqrt{T\log(1/\delta)})$ with probability at least $1-\delta$. This completes the proof of the theorem. ∎
Appendix D Concentration results
Corollary 3.
For a given $k,h,s,a\in[K]\times[H]\times\mathcal{S}\times\mathcal{A}$ (let $n=N_{k,h}(s,a)$), with probability $1-\delta$, it holds,
$|\tilde{Q}_{k,h}(s,a)-\widehat{Q}_{k,h}(s,a)|$ $\leq$ $\sqrt{2\sigma(N_{k,h}(s_{k,h},a_{k,h}))^{2}\log(1/\delta)}.$ (31)
Proof.
The result directly follows form Lemma 17. ∎
Corollary 4.
For some given $k,h,s,a\in[K]\times[H]\times\mathcal{S}\times\mathcal{A}$, the following holds with probability at least $1-\delta$ (with $n=N_{k,h}(s,a)$),
$|\sum_{i=1}^{n}\alpha_{n}^{i}\mathinner{\left(r_{k_{i},h}-R(s,a)+V^{*}_{h+1}(s_{k_{i},h+1})-P_{s,a}V^{*}_{h+1}\right)}|$ $\leq$ $4\sqrt{\frac{H^{3}\log(1/\delta)}{n+1}}.$
Proof.
Let $k_{i}$ denote the index of the episode when $(s,a)$ was visited for the $i_{\text{th}}$ time at step $h$. Set $x_{i}=\alpha_{n}^{i}(r_{k_{i},h}-R_{h}(s,a)+V(s_{i+1})-P_{s_{i},a_{i}}\cdot V)$ and consider filtration $\mathcal{F}_{i}$ as the $\sigma$-field generated by all random variables in the history set $\mathcal{H}_{k_{i},h}$. $r_{k_{i},h}-R_{h}(s,a)+V(s_{i+1})-P_{s_{i},a_{i}}\cdot(V)\leq H+1$. Using the definition of the learning rate (Lemma 16 (b)), we have $\sum_{i}^{n}x^{2}_{i}\leq H(H+1)^{2}/n$. We apply Azuma-Hoeffding inequality (see Lemma 14) combined with a union bound over all $(s,a,h)\in\mathcal{S}\times\mathcal{A}\times[H]$ and all possible values of $n\leq K$ to get the following with probability at least $1-\delta$,
$|\sum_{i=1}^{n}x_{i}|$ $\leq$ $2\sqrt{\frac{2H^{3}\log(1/\delta)}{n}}.$
We complete the proof using the observation $\frac{1}{n+1}\geq\frac{1}{2n},\,n\geq 1$. ∎
Lemma 10.
with probability at least $1-\delta$, the following holds
$\sum_{k=1}^{K}\sum_{h=1}^{H}P_{s_{k,h},a_{k,h}}\cdot(V^{*}_{h+1}-V^{\pi_{k}}_{h+1})-(V^{*}_{h+1}(s_{k,h+1})-V^{\pi_{k}}_{h+1}(s_{k,h+1}))\leq H^{2}\sqrt{2T\log(1/\delta)}$
Proof.
For $x_{k,h}=P_{s_{k,h},a_{k,h}}\cdot(V^{*}_{h+1}-V^{\pi_{k}}_{h+1})-(V^{*}_{h+1}(s_{k,h+1})-V^{\pi_{k}}_{h+1}(s_{k,h+1}))$ and filtration set $\mathcal{H}_{k,h}$ where $k$ is the episode index, $\{x_{k,h},\mathcal{H}_{k,h}\}$ forms a martingale difference sequence with $|x_{k,h}|\leq H$. We complete the proof using Lemma 14 and a union bound. ∎
Lemma 11.
Let $\mathcal{D}_{k,h}$ be the distribution of actions at time step $k,h$ conditioned on the history at the start of step $h$ of the $k_{\text{th}}$ episode, then with probability at least $1-\delta$, for some $h$
$\sum_{k=1}^{K}\mathbb{E}_{a\sim\mathcal{D}_{k,h}}\mathinner{\left[\frac{1}{\sqrt{n_{k,h}(s_{k,h},a)+1}}\right]}-\frac{1}{\sqrt{n_{k,h}(s_{k,h},a_{k,h})+1}}\leq O\mathinner{\left(\sqrt{SA\log(K)}\log(1/\delta)\right)}$
Proof.
For brevity of exposition, let $\mathbb{E}[Z_{k}]\leftarrow\mathbb{E}_{a\sim\mathcal{D}_{k,h}}\mathinner{\left[\frac{1}{\sqrt{n_{k,h}(s_{k,h},a)+1}}\right]}$, $Z_{k}\leftarrow\frac{1}{\sqrt{n_{k,h}(s_{k,h},a_{k,h})+1}}$, and $\mathcal{F}_{k}\leftarrow\mathcal{H}_{k,h}$. Consider,
$\sum_{k=1}^{K}(\mathbb{E}[Z_{k}]-Z_{k})^{2}$ $\leq$ $(\mathbb{E}[Z_{k}])^{2}+Z_{k}^{2}$
(By Jensen’s inequality for $f(x)=x^{2}$)
$\leq$ $\sum_{k=1}^{K}Z_{k}^{2}+\sum_{k=1}^{K}\mathbb{E}[Z_{k}^{2}]$
(by linearity of expectation)
$=$ $\sum_{k=1}^{K}Z_{k}^{2}+\mathbb{E}\mathinner{\left[\sum_{k=1}^{K}Z_{k}^{2}\right]}$
$\leq$ $\sum_{s,a}\sum_{j=1}^{K}\frac{1}{j+1}+\mathbb{E}\mathinner{\left[\sum_{s,a}\sum_{j=1}^{K}\frac{1}{j+1}\right]}$
$\leq$ $2SA\log(K).$
To bound $\sum_{k=1}^{K}(\mathbb{E}[Z_{k}]-Z_{k})$, we apply Bernstein inequality for martingale, Lemma 15, with $K=1$, $d=2SA\log(K)$ to get the required result.
Lemma 12.
$\sum_{k=1}^{K}\frac{1}{\sqrt{N_{k,h}(s_{k,h},a_{k,h})+1}}\leq\mathrm{O}\mathinner{\left(\sqrt{SAK}\right)}$
Proof.
$\sum_{k=1}^{K}\frac{1}{\sqrt{N_{k,h}(s_{k,h},a_{k,h})+1}}$ $\leq$ $\sum_{k=1}^{K}\frac{\sqrt{2}}{\sqrt{N_{k,h}(s_{k,h},a_{k,h})}}$
$\leq$ $\mathrm{O}(1)\sum_{s,a}\sum_{k=1}^{N_{K,h(s,a)}}\sqrt{\frac{1}{k}}$
$\leq$ $\mathrm{O}\mathinner{\left(\sqrt{SAK}\right)}.$
Corollary 5.
The following holds,
$\sum_{h=1}^{H}\sum_{k=1}^{K}\sqrt{4\sigma(N_{k,h}(s_{k,h},a_{k,h}))^{2}\eta}\leq\mathrm{O}\mathinner{\left(\sqrt{H^{4}SAT\eta}\right)}.$
Proof.
From Lemma 12 and (10), we get the result. ∎
Corollary 6.
Let $\mathcal{D}_{k,h}$ be the distribution of actions at time step $k,h$ conditioned on the history at the start of step $h$ of the $k_{\text{th}}$ episode, then with probability at least $1-\delta$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sum_{k=1}^{K}\frac{4}{p_{2}}\mathbb{E}_{a\sim\mathcal{D}_{k,h}}\mathinner{\left[\sqrt{\sigma(N_{k,h}(s_{k,h},a))^{2}}\log(JKH/\delta)\right]}+\sqrt{2\sigma(N_{k,h}(s_{k,h},a_{k,h}))^{2}}\log(JKH/\delta)}$
$\leq$ $O\mathinner{\left(\sqrt{H^{2}SAT}\chi\right)},$
$\chi=\log(JSAT/\delta)$.
Proof.
Using Lemma 11, we have with probability $1-\delta$,
$\sum_{k=1}^{K}\mathbb{E}_{a\sim\mathcal{D}_{k,h}}\mathinner{\left[\frac{1}{\sqrt{n_{k,h}(s_{k,h},a)+1}}\right]}-\frac{1}{\sqrt{n_{k,h}(s_{k,h},a_{k,h})+1}}\leq O\mathinner{\left(\sqrt{SA\log(K)}\log(1/\delta)\right)}$
From Lemma 12 and (10),
$\sum_{k=1}^{K}\frac{4}{p_{2}}\sqrt{\sigma(N_{k,h}(s_{k,h},a))^{2}}\log(JKH/\delta)\leq O\mathinner{\left(\sqrt{H^{2}SAT}\chi\right)},$
which dominates the remaining terms. ∎
Appendix E Technical Preliminaries
See 2
Proof.
For the purpose of this proof, a symmetric sample $\tilde{X}^{\text{alt}}$ also drawn from distribution $\mathcal{D}$ but independent of $\tilde{X}$. Let $\mathcal{O}^{\text{alt}}$ denotes the event when $\tilde{X}^{\text{alt}}\geq X^{*}$ (occurring with probability $p$).
Consider (using notation $\mathbb{E}[\cdot]\leftarrow\mathbb{E}_{\mathcal{D}}[\cdot]$)
$X^{*}-\tilde{X}\leq$ $\mathbb{E}\mathinner{\left[\tilde{X}^{\text{alt}}\mid\mathcal{O}^{\text{alt}}\right]}-\tilde{X}\leq$ $x`x`\mathbb{E}\mathinner{\left[\tilde{X}^{\text{alt}}-\underline{X}\mid\mathcal{O}^{\text{alt}}\right]},$ (32)
where the last inequality holds with probability $1-\delta$ by definition of $\underline{X}$. The law of total expectation suggests,
$\mathbb{E}\mathinner{\left[\tilde{X}^{\text{alt}}-\underline{X}\right]}$ $=$ $\text{Pr}(\mathcal{O}^{\text{alt}})\mathbb{E}\mathinner{\left[\tilde{X}^{\text{alt}}-\underline{X}\mid\mathcal{O}^{\text{alt}}\right]}+\text{Pr}(\overline{\mathcal{O}}^{\text{alt}})\mathbb{E}\mathinner{\left[\tilde{X}^{\text{alt}}-\underline{X}\mid\overline{\mathcal{O}}^{\text{alt}}\right]},$
where $\overline{\mathcal{O}}^{\text{alt}}$ is the compliment of the event $\mathcal{O}^{\text{alt}}$. Now, $\mathbb{E}\mathinner{\left[\tilde{X}^{\text{alt}}\mid\overline{\mathcal{O}}^{\text{alt}}\right]}=\mathbb{E}\mathinner{\left[\tilde{X}^{\text{alt}}\mid\tilde{X}^{\text{alt}}<X^{*}\right]}\leq\mathbb{E}[\tilde{X}^{\text{alt}}]=\mathbb{E}[\tilde{X}]\leq\underline{X}$, where the last inequality is by the assumption made in the lemma. Therefore, the second term in the above is non-negative, and we have
$\mathbb{E}\mathinner{\left[\tilde{X}^{\text{alt}}-\underline{X}\right]}$ $\leq$ $\text{Pr}(\mathcal{O}^{\text{alt}})\mathbb{E}\mathinner{\left[\tilde{X}^{\text{alt}}-\underline{X}\mid\mathcal{O}^{\text{alt}}\right]}.$ (33)
Using $\frac{1}{\text{Pr}(\mathcal{O}^{\text{alt}})}\leq\frac{1}{p}$, $\mathbb{E}[\tilde{X}^{\text{alt}}]=\mathbb{E}[\tilde{X}$], and a union bound to combine (32) and (33), we complete the proof. ∎
Lemma 13.
Let $q^{(1)},q^{(2)},\ldots q^{(M)}$ be $M$ i.i.d. samples such that for any $i$, $q^{(i)}\geq V^{*}$ with probability $p$. Then with probability at least $1-\delta$,
$\max_{i\in M}q^{(i)}\geq V^{*},$
when $M$ is at least $\frac{\log(1/\delta)}{\log(1/(1-p))}$.
Proof.
For a given index $i$, the probability that $q^{(i)}<V^{*}$ is at most 1p1-p. Therefore, by independence of samples, the probability of $\max_{i\in J}q^{(i)}<V^{*}$ is at most $(1-p)^{M}$. Therefore, the lemma statement follows by setting $M=\frac{\log(1/\delta)}{\log(1/(1-p))}$. ∎
Lemma 14 (Corollary 2.1 in Wainwright [2019]).
Let $(\{A_{i},\mathcal{F}_{i}\}_{i=1}^{\infty})$ be a martingale difference sequence, and suppose $|A_{i}|\leq d_{i}$ almost surely for all $i\geq 1$. Then for all $\eta\geq 0$,
$\mathbb{P}\left[|\sum_{i=1}^{n}A_{i}|\geq\eta\right]\leq 2\exp\mathinner{\left(\frac{-2\eta^{2}}{\sum_{i=1}^{n}d_{i}^{2}}\right)}.$ (34)
In other words, with probability at most $\delta$, we have,
$|\sum_{i=1}^{n}A_{i}|\geq\sqrt{\frac{ln\mathinner{\left(2/\delta\right)}\sum_{i=1}^{n}d_{i}^{2}}{2}}$ (35)
Lemma 15 (Lemma A8 in Cesa-Bianchi and Lugosi [2006]).
Let $(\{A_{i},\mathcal{F}_{i}\}_{i=1}^{\infty})$ be a martingale difference sequence, and suppose $|A_{i}|\leq K$ almost surely for all $i\geq 1$. Let $S_{i}=\sum_{j=1}^{i}A_{i}$ be the associated martingale. Denote the sum of the conditional variances by
$v_{n}^{2}=\sum_{t=1}^{n}\mathbb{E}[A_{i}^{2}\mid\mathcal{F}_{1-1}].$
Then for all constants $t,d>0$,
$\mathbb{P}\mathinner{\left[\max_{i}S_{i}\geq t\,\&\,,v_{n}^{2}\leq d\right]}\leq\exp\mathinner{\left(-\frac{t^{2}}{2(d+Kt/3)}\right)}.$
Lemma 16 (Lemma 4.1 in Jin et al. [2018]).
The following holds:
  1. (a)
    $\frac{1}{\sqrt{n}}\leq\sum_{i=1}^{n}\frac{\alpha^{i}_{n}}{\sqrt{i}}\leq\frac{2}{\sqrt{n}}$.
  2. (b)
    $\max_{i\in n}\alpha_{n}^{i}\leq\frac{2C}{t}$ and $\sum_{i=1}^{n}(\alpha_{n}^{i})^{2}\leq\frac{2C}{t}.$
  3. (c)
    $\sum_{n=i}^{\infty}\alpha^{i}_{n}\leq 1+1/C$.
Lemma 17 (Gaussian tail bound).
For a Gaussian random variable $X\sim\mathcal{N}(\mu,\sigma^{2})$, it follows with probability at least $1-\delta$,
$\text{Pr}\mathinner{\left(\mathinner{\!\left\lvert X-\mu\right\rvert}\leq\sigma\sqrt{2\log(1/\delta)}\right)}$
Proof.
The proof follows by instantiating Chernoff-style bounds for the given Gaussian random variable. ∎
Appendix F Sharper regret using Bernstein concentration
1:Input: Parameters: $\delta\in(0,1)$. Set $J\mathrel{\mathop{\ordinarycolon}}=J(\delta)$.
2:Initialize: $\widehat{Q}_{H+1}(s,a)=\widehat{V}_{H+1}(s)=0,\,\forall(s,a)\in\mathcal{S}\times\mathcal{A}$, and $\widehat{Q}_{h}(s,a)=\widehat{V}_{h}(s)=H\,\&\,N_{h}(s,a)=0,N_{h}^{\text{tar}}(s)=0,\mu_{h}(s,a)=\gamma_{h}(s,a)=0\,\forall(s,a,h)\in\mathcal{S}\times\mathcal{A}\times[H]$.
3:for episodes $k=1,2,\ldots$ do
4:  Observe $s_{1}$.
5:  for step $h=1,2,\ldots,H$ do
6:    /* Play arg max action of the sample of $Q_{h}$ */
7:   Sample $\forall a\in\mathcal{A}$ $\tilde{Q}_{h}(s_{h},a)\sim\mathcal{N}(\widehat{Q}_{h}(s_{h},a),v_{b}(N_{h}(s_{h},a),h,s_{h},a))$.
8:   Play $a_{h}=\operatorname*{arg\,max}_{a_{\mathcal{A}}}\tilde{Q}_{h}(s_{h},a)$.
9:    Use observations to construct one step lookahead target $z$ */
10:   Observe $r_{h}$ and $s_{h+1}$.
11:   $z\leftarrow\texttt{ConstructTarget}(r_{h},s_{h+1},\widehat{Q}_{h+1},N_{h+1})$.
12:    /*Use the observed reward and next state to update $Q_{h}$ distribution */
13:   $n\mathrel{\mathop{\ordinarycolon}}=N_{h}(s_{h},a_{h})\leftarrow N_{h}(s_{h},a_{h})+1$.
14:   $\widehat{Q}_{h}(s_{h},a_{h})\leftarrow(1-\alpha_{n})\widehat{Q}_{h}(s_{h},a_{h})+\alpha_{n}z$.
15:   $\mu_{h}(s_{h},a_{h})\leftarrow\mu_{h}(s_{h},a_{h})+(z-r_{h})$.
16:   $\gamma_{h}(s_{h},a_{h})\leftarrow\gamma_{h}(s_{h},a_{h})+(z-r_{h})^{2}$.
17:   Calculate $b_{h+1}(s_{h},a_{h})\leftarrow(\gamma_{h}(s_{h},a_{h})-\mu_{h}(s_{h},a_{h})^{2})/n$.
18:   $\sqrt{v_{b}(n,h,s_{h},a_{h})}\leftarrow\min\{c(\sqrt{\frac{H}{n+1}\cdot(b_{h+1}(s_{h},a_{h})+H)\eta}+\frac{\sqrt{H^{7}SA\eta}\chi}{n+1}),\sqrt{64\frac{H^{3}}{n+1}}\}$.
19:  end for
20:end for
Algorithm 3 Randomized Q-learning
In this section, we provide a sketch of the extension of Algorithm 1 to a randomized Q-learning procedure that uses Bernstein concentration based variance. This extension closely follows that in Jin et al. [2018] using some of the techniques developed in the proof of Theorem 2.
We want to account for the variance in the transitions. To this end, we define some additional notations. The variance in transition for any $(s,a)$ is defined using the variance operator $\mathbb{V}_{h}$ as below,
$[\mathbb{V}_{h}V_{h+1}]_{s,a}\coloneqq\mathbb{E}_{s^{\prime}\sim P_{h,s,a}}\mathinner{\left[V_{h+1}(s^{\prime})-[P_{h,s,a}V_{h+1}]_{s,a}\right]}^{2}.$ (36)
The empirical variance for any $(s,a)$ for any $n\leftarrow N_{k,h}(s,a)$ is given by,
$\widehat{\mathbb{V}}_{n}\overline{V}_{h+1}(s,a)=\frac{1}{n}\sum_{i=1}^{n}\left[\overline{V}_{k_{i},h+1}(s_{k_{i},h+1})-\frac{1}{n}\sum_{i=1}^{n}\overline{V}_{k_{i},h+1}(s_{k_{i},h+1})\right]^{2}$ (37)
$\sqrt{v_{b}(n,h,s,a)}\coloneqq\min\mathinner{\left\{c(\sqrt{\frac{H}{n+1}\cdot(\widehat{\mathbb{V}}_{n}\overline{V}_{h+1}(s,a)+H)\eta}+\frac{\sqrt{H^{7}SA\eta}\cdot\tau}{n+1}),\sqrt{64\frac{H^{3}}{n+1}}\right\}}.$ (38)
Theorem 3.
The cumulative regret of Algorithm 3 in $K$ episode satisfies with probability at least $1-\delta$,
$\sum_{k=1}V^{*}_{1}(s_{k,1})-\sum_{k=1}^{K}\sum_{h=1}^{H}R_{h}(s_{k,h},a_{k,h})\leq O\mathinner{\left(\sqrt{H^{3}SAT\eta}\chi\right)},$
where $\chi=\log(JSAT/\delta)$ and $\eta=\log(SAKH/\delta)$.
F.1 Proof of Theorem 3
The main mathematical reasoning closely follows that in the proof of Theorem 2 of Jin et al. [2018] with specific differences arising due to constant probability optimism and the definition of $\overline{V}_{k,h}$. For any $k,h,s,a$ with $n=N_{k,h}(s,a)$, we have $v_{b}(n,h,s,a)\leq 64\frac{H^{3}}{n+1}$. Therefore, Corollary 2 and Lemma 6 apply as they are. Hence, we get with probability at least $1-\delta$ for all $h$
$\sum_{k=1}^{K}r_{k,h}\coloneqq\sum_{k=1}^{K}V^{*}_{h}(s_{k,h})-V^{\pi_{k}}_{h}(s_{k,h})$ $\leq$ $O\mathinner{\left(\sqrt{H^{4}SAT}\chi\right)},$ (39)
where $\chi=\log(JSAT/\delta)$. Further, following the steps and notations of the proof of Lemma 6 (see (28), we have with probability at least $1-\delta$,
$\sum_{k=1}^{K}\tilde{Q}_{k,1}(s_{k,1},a_{k,1})-V^{\pi_{k}}_{1}(s_{k,1})$ $\leq$ $eSAH^{2}+e\sum_{h=1}^{H}\sum_{k=1}^{K}\mathinner{\left(\sqrt{4v_{b}(n,h,s_{k,h},a_{k,h})\eta}+m_{k,h}\right)}$ (40)
$+\frac{e}{p_{2}}\sum_{h=1}^{H}\sum_{k=1}^{K}F({k,h+1},\delta/KH)).$
Due to our observation that $v_{b}(n,h,s_{k,h},a_{k,h})\leq 64\frac{H^{3}}{n+1}$, Lemma 11 and Corollary 6 hold, therefore we get from (40),
$\sum_{k=1}^{K}\tilde{Q}_{k,1}(s_{k,1},a_{k,1})-V^{\pi_{k}}_{1}(s_{k,1})$ $\leq$ $eSAH^{2}+\sum_{h=1}^{H}\sum_{k=1}^{K}O\mathinner{\left(\sqrt{v_{b}(n,h,s_{k,h},a_{k,h})}\chi+m_{k,h}\right)},$
where $\chi=\log(JSAT/\delta)$. We wish to bound
$\sum_{k=1}^{K}\sum_{h=1}^{H}v_{b}(n,h,s_{k,h},a_{k,h})$ (41)
$\leq$ $\sum_{k=1}^{K}\sum_{h=1}^{H}c(\sqrt{\frac{H}{n+1}\cdot(\widehat{\mathbb{V}}_{n}\overline{V}_{h+1}(s_{k,h},a_{k,h})+H)\eta}+\frac{\sqrt{H^{7}SA\eta}\cdot\tau}{n+1}),$ (42)
Consider,
$\sum_{k=1}^{K}\sum_{h=1}^{H}\frac{\sqrt{H^{7}SA\eta}\chi}{N_{k,h}(s_{k,h},a_{k,h})+1}\leq O(\sqrt{H^{9}S^{3}A^{3}\chi^{5}}).$ (43)
Further,
$\sum_{k=1}^{K}\sum_{h=1}^{H}\sqrt{\frac{H}{n+1}\cdot(\widehat{\mathbb{V}}_{n}\overline{V}_{k,h+1}(s_{k,h},a_{k,h})+H)}$
$\leq$ $\sum_{k=1}^{K}\sum_{h=1}^{H}\sqrt{\frac{H}{n+1}\cdot\widehat{\mathbb{V}}_{n}\overline{V}_{k,h+1}(s_{k,h},a_{k,h})}+\sum_{k=1}^{K}\sum_{h=1}^{H}\sqrt{\frac{H^{2}}{n+1}}$
$\leq$ $\sqrt{\sum_{k=1}^{K}\sum_{h=1}^{H}\widehat{\mathbb{V}}_{n}\overline{V}_{k,h+1}(s_{k,h},a_{k,h})H}+\sqrt{H^{3}SAT\eta},$ (44)
where the last inequality follows from Lemma 12. Since we have Lemma 22, we can follow the steps in (C.16) of Jin et al. [2018] to get
$\sum_{k=1}^{K}\sum_{h=1}^{H}\widehat{\mathbb{V}}_{n}\overline{V}_{k,h+1}(s_{k,h},a_{k,h})\leq O(HT).$
This gives us
$\sum_{k=1}^{K}\sum_{h=1}^{H}v_{b}(n,h,s_{k,h},a_{k,h})\leq O(\sqrt{H^{3}SAT\eta}+\sqrt{H^{9}S^{3}A^{3}\chi^{5}})$
Thus we have,
$\sum_{k=1}^{K}\tilde{Q}_{k,1}(s_{k,1},a_{k,1})-V^{\pi_{k}}_{1}(s_{k,1})\leq O(\sqrt{H^{3}SAT\eta}\chi+\sqrt{H^{9}S^{3}A^{3}\chi^{5}})$
Finally, combining with Corollary 2, we complete the proof.
F.2 Supporting lemma
Corollary 7 (Corollary of Lemma 4).
We have for all $s,a,h,k\in\mathcal{S}\times\mathcal{A}\times[H]\times[K]$ with probability at least $1-\delta$,
$\widehat{Q}_{k,h}(s,a)-Q^{*}_{h}(s,a)\leq O\mathinner{\left(\sqrt{\frac{H^{3}\eta}{n+1}}\right)}+\alpha_{n}^{0}H+\sum_{i=1}^{n}\alpha_{n}^{i}\mathinner{\left(\overline{V}_{k_{i},h+1}(s_{k_{i},h+1})-V^{*}_{h+1}(s_{k_{i},h+1})\right)},$ (45)
where $n=N_{k,h}(s,a)$, and $\eta=\log(SAKH/\delta)$.
Proof.
The proof follows from using (38) in place of $\sigma(N_{k,h}(s,a))^{2}$ in Lemma 4. ∎
Lemma 18 (Based on Lemma C.7 in  Jin et al. [2018]).
Suppose (45) in Corollary 7 holds. For any $h\in[H]$, let $\beta_{k,h}\coloneqq\overline{V}_{k,h}(s_{k,h})-V^{*}_{h}(s_{k,h})$ and let $w=(w_{1},\ldots,w_{k})$ be non-negative weight vectors, then we have with probability at least $1-\delta$,
[…]
where $\chi=\log(JSAT/\delta)$.
Proof.
For any fixed $k,h$, let $n\leftarrow N_{k,h}(s_{k,h},a_{k,h})$ and $\phi_{k,h}=\tilde{Q}_{k,h}(s_{k,h},a_{k,h})-V^{*}_{h}(s_{k,h})$. Then we have
$\beta_{k,h}$ $=$ $\overline{V}_{k,h}(s_{k,h})-V^{*}_{h}(s_{k,h})$ (46)
(from Lemma 9 with probability at least $1-2\delta$)
$\leq$ $\phi_{k,h}+\frac{1}{p2}F(k,h,\delta)$
(from Lemma C.5 with probability $1-\delta$)
$\leq$ $\widehat{Q}_{k,h}(s_{k,h},a_{k,h})-Q^{*}_{h}(s_{k,h},a_{k,h})+\sqrt{2v(n)\eta}+\frac{1}{p_{2}}F(k,h,\delta)$
(from Corollary 7 with probability $1-\delta$)
$\leq$ $\alpha_{n}^{0}H+O\mathinner{\left(\sqrt{\frac{H^{3}\eta}{n+1}}\right)}+\sum_{i=1}^{n}\alpha_{n}^{i}\beta_{k_{i},h+1}+\frac{1}{p_{2}}F(k,h,\delta)$
We now compute the summation $\sum_{k=1}^{K}w_{k}\beta_{k,h}$. We follow the proof of Lemma C.7 in Jin et al. [2018], with the only difference being the term $\sum_{k=1}^{K}\frac{w_{k}}{p_{2}}F(k,h,\delta)$, which we bound below. From the proof of Corollary 6.
[…]
where $\chi=\log(JSAT/\delta)$. Other terms are evaluated in the same way as the proof of Lemma C.7 in Jin et al. [2018]. ∎
Lemma 19 (Based on Lemma C.3 of Jin et al. [2018]).
For any episode $k\in[K]$ with probability $1-\delta/K$, if Corollary 7 holds for all $k^{\prime}<k$, the for all $s,a,h\in\mathcal{S}\times\mathcal{A}\times[H]$:
$\mathinner{\!\left\lvert[\mathbb{V}_{h}V_{h+1}]_{s,a}-\widehat{\mathbb{V}}_{n}\overline{V}_{h+1}(s,a)\right\rvert}\leq O\mathinner{\left(\frac{SA\sqrt{H^{9}}\chi^{2}}{n}+\sqrt{\frac{H^{7}SA\chi^{2}}{n}}\right)},$
where $n=N_{k,h}(s_{k,h},a_{k,h})$, $\chi=\log(JSAT/\delta)$.
Proof.
The proof is almost identical to that of Lemma C.3 of Jin et al. [2018] except we Lemma 18 instead of Lemma C.7 of Jin et al. [2018]. ∎
Lemma 20.
(Bernstein concentration) For some given $k,h,s,a\in[K]\times[H]\times\mathcal{S}\times\mathcal{A}$, the following holds with probability at least $1-\delta$ (with $n=N_{k,h}(s,a)$),
$|\sum_{i=1}^{n}\alpha_{n}^{i}\mathinner{\left(r_{k_{i},h}-R(s,a)+V^{*}_{h+1}(s_{k_{i},h+1})-P_{s,a}V^{*}_{h+1}\right)}|$ $\leq$ $\sqrt{v_{b}(n,h,s,a)},$
where
$v_{b}(n,h,s,a)\coloneqq\min\mathinner{\left\{c(\sqrt{\frac{H}{n+1}\cdot(\widehat{\mathbb{V}}_{n}\overline{V}_{h+1}(s,a)+H)\eta}+\frac{\sqrt{H^{7}SA\eta}\cdot\tau}{n+1}),64\frac{H^{3}}{n+1}\right\}}.$
Proof.
Let $k_{i}$ denote the index of the episode when $(s,a)$ was visited for the $i_{\text{th}}$ time at step $h$. Set $x_{i}=\alpha_{n}^{i}(r_{k_{i},h}-R_{h}(s,a))$, $y_{i}=\alpha_{n}^{i}(V(s_{k_{i+1}})-P_{s_{i},a_{i}}\cdot V)$ and consider filtration $\mathcal{F}_{i}$ as the $\sigma$-field generated by all random variables in the history set $\mathcal{H}_{k_{i},h}$. We apply Azuma-Hoeffding (Lemma 14) to calculate $|\sum_{i=1}^{n}x_{i}|$ and Azuma-Bernstein for $|\sum_{i=1}^{n}y_{i}|$. Consider with probability $1-\delta$,
$\mathinner{\!\left\lvert\sum_{i=1}^{n}\alpha_{n}^{i}(V(s_{k_{i+1}})-P_{s_{i},a_{i}}\cdot V)\right\rvert} \displaystyle\leq$ $O(1)\cdot\mathinner{\left[\sum_{i=1}^{n}\sqrt{(\alpha_{n}^{i})^{2}[\mathbb{V}_{h}V^{*}_{h+1}]_{s,a}\eta}+\frac{H^{2}\eta}{n}\right]}$
$\leq$ $O(1)\cdot\mathinner{\left[\sum_{i=1}^{n}\sqrt{\frac{H}{n}[\mathbb{V}_{h}V^{*}_{h+1}]_{s,a}\eta}+\frac{H^{2}\eta}{n}\right]}.$
Using Lemma 19 and a suitable union bound, we have with probability $1-\delta$,
$\mathinner{\!\left\lvert\sum_{i=1}^{n}\alpha_{n}^{i}(V(s_{k_{i+1}})-P_{s_{i},a_{i}}\cdot V)\right\rvert} \displaystyle\leq$ $O(1)\cdot\mathinner{\left[\sum_{i=1}^{n}\sqrt{\frac{H}{n+1}(\widehat{\mathbb{V}}_{n}\overline{V}_{h+1}(s,a)+H)\eta}+\frac{\sqrt{H^{7}SA\eta}\cdot\chi}{n+1}\right]}$
$\leq$ $\sqrt{v_{b}(n,h,s,a)}.$
where the last term dominates the concentration of $|\sum_{i=1}^{n}x_{i}|$. ∎
Lemma 21 (Based on Lemma 8).
The samples from the posterior distributions and the mean of the posterior distributions as defined in Algorithm 3 satisfy the following properties: for any episode $k\in[K]$ and index $h\in[H]$,
  1. (a)
    (Posterior distribution mean) For any given $s,a$, with probability at least $\Phi(-1)-\frac{\delta}{H}-\delta$,
    $\widehat{Q}_{k,h}(s,a)\geq Q^{*}_{h}(s,a)-\sqrt{v_{b}(n,h,s,a)}.$ (47)
  2. (b)
    (Posterior distribution sample) For any given $s,a$, with probability at least $p_{1}$ ($p_{1}=\Phi(-1)-\frac{\delta}{H}-\delta$),
    $\tilde{Q}_{k,h}(s,a)\geq Q^{*}_{h}(s,a).$ (48)
  3. (c)
    (In Algorithm 2) With probability at least $1-\frac{\delta}{KH}$,
    $\overline{V}_{k,h}(s_{k,h})\geq V^{*}_{h}(s_{k,h}).$ (49)
Here $\log(1/\delta^{\prime})=4$ and $\delta$ is a parameter of the algorithm used to define the number of samples $J$ used to compute the target $\overline{V}$.
Proof.
The proof is identical to that of Lemma 8 except that we use Lemma 20 instead of Corollary 4. ∎
Lemma 22 (Based on Lemma C.6 of Jin et al. [2018]).
With probability at least $1-4\delta$, we have the following for all $k,h\in[K]\times[H]$,
$\widehat{\mathbb{V}}_{n}\overline{V}_{k,h}(s_{k,h},a_{k,h})-\mathbb{V}_{h}V^{\pi_{k}}_{h+1}(s_{k,h},a_{k,h})$ $\leq$ $2HP_{s_{k,h},a_{k,h}}\cdot(V^{*}_{h+1}-V^{\pi_{k}}_{h+1})$
$+O\mathinner{\left(\frac{SA\sqrt{H^{9}\chi^{4}}}{n}+\sqrt{\frac{H^{7}SA\chi^{2}}{n}}\right)},$
where $n=N_{k,h}(s_{k,h},a_{k,h})$ and $\chi=\log(JSAT/\delta)$.
Proof.
Consider,
$\widehat{\mathbb{V}}_{n}\overline{V}_{k,h}(s_{k,h},a_{k,h})-\mathbb{V}_{h}V^{\pi_{k}}_{h+1}(s_{k,h},a_{k,h})$ $\leq$ $\mathinner{\!\left\lvert\widehat{\mathbb{V}}_{n}\overline{V}_{k,h}(s_{k,h},a_{k,h})-\mathbb{V}_{h}V^{*}_{h+1}(s_{k,h},a_{k,h})\right\rvert}$
$+\mathinner{\!\left\lvert\mathbb{V}_{n}V^{*}_{h+1}(s_{k,h},a_{k,h})-\mathbb{V}_{h}V^{\pi_{k}}_{h+1}(s_{k,h},a_{k,h})\right\rvert}$,
where for the first term we apply Lemma 19 (which holds when Corollary 7 and Lemma 18 hold) and for the second term we have from the definition of variance,
$\mathinner{\!\left\lvert\mathbb{V}_{n}V^{*}_{h+1}(s_{k,h},a_{k,h})-\mathbb{V}_{h}V^{\pi_{k}}_{h+1}(s_{k,h},a_{k,h})\right\rvert} \displaystyle\leq$ $2HP_{s_{k,h},a_{k,h}}\cdot(V^{*}_{h+1}-V^{\pi_{k}}_{h+1}).$